### v1

# Best approximation of a three-variable function by sum of one-variable coordinate functions

Abstract
Consider the following approximation problem of a continuous function of three variables by the sum of three continuous functions of one variable:
$E(f,\Omega)=\inf||f(x,y,z)-\phi(x)-\psi(y)-\omega(z)||_{\infty}=?$
where $f(x,y,z)$ is a given continuous function defined on $\Omega=[0,1]^{3}$ and the infimum runs over all triplets of continuous functions $\phi(x),\psi(y),\omega(z)$ defined on the unit interval $[0,1]$. In this paper, we will prove a formula to calculate the error $E(f,\Omega)$ under certain conditions.
1,2,3 Baku State University, Baku, Azerbaijan
1Institute of Mathematics and Mechanics, Baku, Azerbaijan
1Center for Mathematics and its Applications, Khazar University, Baku, Azerbaijan
e-mail: 1aliyevrashid@mail.ru, 2 quliyevvuqar.213@gmail.com, 3 amilcbiyev23@gmail.com
1 Introduction
1.1 Problem statement in a general form and previous results
Let $f:X\to\mathbb{R}$ be a continuous function in the $n$ variables, where $X=X_{1}\times X_{2}\times\dots X_{n}$ is the Cartesian product of the compact Hausdorff spaces $n$. Consider the approximation of $f$ by means of $g_{i},i=1,\dots,n$, where $g_{i}$ is a continuous function on $X_{i}$ for $i=1,\dots,n$. In other words, the problem asks for
$E(f,X)=\inf_{(g_{1},g_{2},...,g_{n})}||f(x_{1},x_{2},\dots,x_{n})-\sum_{i=1}^{n}g_{i}(x_{i})||_{C(X)}=?$ (1.1)
where the infimum runs over all $n$-tuples of functions $(g_{1},g_{2},...,g_{n})$ satisfying $g_{i}\in C(X_{i}),i=1,\dots,n$.
In general, this problem is related to Hilbert’s 13th problem. Using A.S. Kronrod’s [16] results, A.N. Kolmogorov [14] showed that any function $f\in C([0,1]^{n})$ is representable as a superposition of continuous functions of three variables. Developing this method, V.I. Arnold [1], [2] proved that each continuous function of three variables can be represented by superposition of continuous functions of two variables. This is a negative answer to Hilbert’s conjecture. Further, A.N. Kolmogorov [15] proved the following remarkable, stronger theorem in a simple, elementary and elegant way:
Kolmogorov’s Theorem. There exist increasing functions $\phi_{pq}\in C([0,1])$, $p=1,...,n$, $q=1,...,2n+1$, such that an arbitrary given function
$f(x_{1},...,x_{n})\in C([0,1]^{n})$
can be represented in the form
$f(x_{1},...,x_{n})=\sum_{q=1}^{2n+1}g_{q}\left(\sum_{p=1}^{n}\phi_{pq}(x_{p})\right),$
where $g_{q}\in C([0,1]$ depend on $f$.
Thus, every continuous function on the $n$ variables can be represented by superposition of continuous functions of one variable and the simplest function of two variables, namely, the sum of variables.
In addition to the problems of finding a precise expression for a given function in terms of a combination of functions of fewer variables, it is natural to consider the best approximation of functions of several variables by combinations of functions of fewer variables. These problems arise in various areas of mathematics, for example, in solving integral equations (see [5]), in solving functional equations (see [4]), in saving information (see [6]), in scaling matrices (see [8], [9]), etc. In space $C([0,1]^{n})$ this problem was first studied by S. P. Diliberto and E. G. Straus [6] (the problem was proposed by RAND Corporation for the most efficient distribution of data in the small memory of the computers of the day). The same problem was investigated by Yu. P. Ofman [20] for bounded bivariant functions, and for more general sets by D. E. Marshall and A. G. O’Farrell in [17]. Diliberto and Straus gave a ’leveling algorithm’ to find the approximation error (1.1) for $n=2$ and propose a theorem that will ’analogously’ hold for $n\geq 3$. However, G. Aumann in his work [3] has shown that while the proof in [6] was complete for case $n=2$, the generalization of the algorithm suggested for $n\geq 3$ was incorrect. A direct counterexample refuting the ”leveling algorithm” in $n\geq 3$ was given by V.A. Medvedev [18]. In the paper [3] G. Aumann showed a necessary and sufficient condition on the class of approximating functions that the ”leveling algorithm” works.
Another generalization approach for $n\geq 3$ to the problem starts with the paper [10] by M. Golomb, where he gave an important duality formula for the problem (1.1). Moreover, the works of G. Aumann and [17] by D.E. Marshall and A.G. O’Farrell have shown that the proof of the duality given in [10] also had a gap. Future, based on the results of S.Ya. Khavinson [13] and K.G. Navada [19], V. Ismailov [12] proved that the formula given by M. Golomb is valid in a stronger form.
Note that for the two-dimensional case, under the condition $f_{xy}\geq 0$ T.J. Rivlin and R.J. Sibher [21] established that
$E(f,[0,1]^{2})=\frac{1}{4}[f(1,1)-f(1,0)-f(0,1)+f(0,0)].$
L. Flatto [7] gave a different proof of this formula in his work, where a generalization to the approximation of a $n$ variable function by $n-1$ variable functions was suggested under similar conditions. In case $n\geq 3$ (even with additional conditions), there is no finite algorithm for calculate the error $E(f,X)$. In this paper, under the conditions $\Delta_{x}\Delta_{y}f\geq 0$, $\Delta_{x}\Delta_{z}f\geq 0$, $\Delta_{y}\Delta_{z}f\geq 0$, we will prove a formula to calculate the error (1.1) for the three-variable continuous function on $[0,1]^{3}$.
1.2 Basic definitions and formulas
Let’s give some definitions that will be used throughout this paper. Most definitions in this subsection are taken from the paper [12] (also can be found in the monograph [11]).
Definition 1.1.
A set of points $p=(x_{1},x_{2},...,x_{m}),x_{i}\in X$ where the set $X$ is the Cartesian product $X=X_{1}\times X_{2}\times\dots X_{n}$, is called a projection cycle if there exist some nonzero real numbers $\lambda_{1},\lambda_{2},..\lambda_{m}$ such that
$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\forall j,0\leq j\leq n\text{ and }\forall\xi\in X_{j}$ (1.2)
where $\pi_{j}(x)$ is the projection of $x$ into $X_{j}$ and $\xi$ is an arbitrary element of the set $X_{j}$.
The definition of a projection cycle goes back to the M.Golomb [10]. The above-given definition is an equivalent form of Golomb’s definition taken from [12].
Remark 1.1.
For $n=3$, the left-hand side of the equation (1.2) becomes the sum over planes parallel to one of the coordinate planes. Thus, if a projection-cycle has a point on a plane parallel to one of the coordinate planes, then it has at least one other point with an opposite sign on the same plane.
Definition 1.2.
A projection cycle $p=(x_{1},x_{2},...,x_{m})$ together with a vector $\lambda=(\lambda_{1},\lambda_{2},..\lambda_{m})$ satisfying the equation (1.2) is called a projection cycle-vector and denoted by $(p,\lambda)$. The number of points of $p$ is called the length of the projection-cycle vector $(p,\lambda)$. And $\lambda_{i}=\lambda(x_{i})$ is called the weight at point $x_{i}$.
When we are given a weak projection cycle $p$ and a nonzero vector $\lambda$ satisfying the equation (1.3), by $(p,\lambda)$ we mean the projection cycle-vector based on the points of $p$ corresponding to nonzero components of $\lambda$.
Definition 1.3.
A projection cycle is called a minimal projection cycle if it has no proper subset that is a projection cycle. A projection cycle-vector $(p,\lambda)$ is called minimal if the set of points of $p$ corresponding to nonzero components of $\lambda$ is a minimal projection cycle.
Definition 1.4.
A projection cycle-vector $(p,\lambda)$ is called orthonormal if
$\sum|\lambda_{i}|=1$
Definition 1.5.
A set of points $p=(x_{1},x_{2},...,x_{m}),x_{i}\in X$ where the set $X$ is the Cartesian product $X=X_{1}\times X_{2}\times\dots X_{n}$, is called a weak projection cycle if there exist some real numbers $\lambda_{1},\lambda_{2},..\lambda_{m}$ such that, at least one $\lambda_{i}$ is nonzero and,
$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\forall j,0\leq j\leq n\text{ and }\forall\xi\in X_{j}$ (1.3)
where $\pi_{j}(x)$ is the projection of $x$ into $X_{j}$ and $\xi$ is an arbitrary element of the set $X_{j}$.
The difference between a projection cycle and a weak projection cycle is that the second may contain points with zero weight. Thus, any projection-cycle is a weak projection cycle, but the converse is not true in general. For example, the set of points $\{(0,0),(0,1),(1,0),(1,1),(2,2)\}$ in $\mathbb{R}^{2}$ is a weak projection cycle but not a projection cycle.
The concept of a minimal projection cycle and a weak projection cycle was defined by name loop and weak loop in [19] by K.G. Navada.
The following theorem expresses Golomb’s formula (proposed in [10]), in revisited and improved form, from [12] (See: Theorem 2.8 in [12] or Theorem 3.37 in [11]).
Theorem 1.1 (Golomb’s Formula).
The error of approximation in (1.1) can be given as
$E(f,X)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (1.4)
where the supremum runs over all minimal projection cycles-vectors on the set $X$.
Remark 1.2.
However, it is not a calculable formula; the formula (1.4) can be improved by comparing orthonormal sums over distinct projection-cycles under conditions like [21] or [7] used in their works to get a formula. In this paper, we will find a formula for the case $n=3$ and $X=[0,1]^{3}$.
Remark 1.3.
For simplicity, we use the term ”projection cycle-vector” instead of ”projection cycle, vector pair” used in [12] and [11].
2 Three variable case
Let $\Omega=[0,1]^{3}$ denote the unit cube and $f:\Omega\rightarrow\mathbb{R}$ be a given continuous function on $3$ variables. Consider the problem:
$E(f,\Omega)=\inf||f(x,y,z)-\phi(x)-\psi(y)-\omega(z)||_{\infty}=?$
where the infimum runs over all triplets of continuous functions $\phi(x),\psi(y),\omega(z)$ defined on the unit interval $[0,1]$.
Let, $\Delta_{(x,y,z)}f$ be defined as
$\Delta_{(x,y,z)}f(u,v,w)=f(u+x,v+y,w+z)-f(u,v,w)$
For simplicity, let’s denote
$\Delta_{x}=\Delta_{(x,0,0)},\Delta_{y}=\Delta_{(0,y,0)},\Delta_{z}=\Delta_{(0,0,z)}$
Assume $f$ satisfies the following conditions at any point of $\Omega$
$\begin{cases}\Delta_{x}\Delta_{y}f\geq 0\\ \Delta_{x}\Delta_{z}f\geq 0\\ \Delta_{y}\Delta_{z}f\geq 0\end{cases}$ (2.1)
If $f$ has mixed derivatives with respect to any pair of variables at any point of $\Omega$,then the inequalities (2.1) are equivalent to
$\begin{cases}f_{xy}\geq 0\\ f_{yz}\geq 0\\ f_{xz}\geq 0\end{cases}$ (2.2)
Let, for any finite sets $S$, the set $X_{S}$ denote the set of all projection cycle-vectors, which, each point belongs to the set $S$ and $X_{S,min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in X_{S}$.
We obtain the following result.
Theorem 2.1.
Let, $f\in C(\Omega)$ be a function satisfying the relations (2.1), then there exist a finite set $U=U_{f}$ such that,
$E(f,\Omega)=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z},min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$ (2.3)
where, $U_{x},U_{y},U_{z}$ are the projections of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively.
Remark 2.1.
For $n=2$, above-mentioned papers [6],[7], [11] gives similar teorems. But, while the set $U$ was the corners of the unit square, which is independent of $f$ in the case $n=2$, $U$ will be dependent on the function $f$ for $n=3$.
Below, we will prove the Theorem $2.1$ and moreover will explicitly construct the set $U$.
We start with applying Golomb’s Formula (1.4) and get:
$E(f,\Omega)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.4)
where, the supremum runs over all minimal projection cycle-vectors $(p,\lambda)$ on the set $\Omega$.
To proceed further, we will search the supremum (2.4) as
$E(f,\Omega)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)$ (2.5)
where the last supremum runs over all triples of finite sets $A,B,C\subset[0,1]$ Obviously, for any fixed $U$ and any projection cycle-vector $(p,\lambda)$, there exist at least one triple of sets $A,B,C\subset[0,1]$ with $U\subset A\times B\times C$ such that $(p,\lambda)\in X_{A\times B\times C}$. On the other hand choosing the sets $A,B,C$ finite ensures that the above maximum over $X_{A\times B\times C}$ is attained. Hence, the relation (2.5) is well-defined.
Before proving the Theorem $2.1$ we have to prove some facts. Later, we will see that to prove the Theorem 2.1 for the class of functions satisfying the conditions (2.2) is equivalent to proving it for functions satisfying the following stricter inequalities
$\begin{cases}\Delta_{x}\Delta_{y}f>0\\ \Delta_{x}\Delta_{z}f>0\\ \Delta_{y}\Delta_{z}f>0\end{cases}$ (2.6)
Thus, below here we will work with functions satisfying (2.6).
2.1 Optimal Projection Cycle-Vectors on a Finite Set
First, we will investigate some properties of projection cycle-vectors that give the following maximum
$\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.7)
where $A,B,C\subset[0,1]$ are arbitrary chosen sets with $\{0,1\}\subset A\cap B\cap C$.
Let’s give some further definitions and auxiliary facts.
Definition 2.1.
Let’s call two points $(\xi_{1},\gamma_{1},\zeta_{1})$ and $(\xi_{2},\gamma_{2},\zeta_{2})$ of the space $\mathbb{R}^{3}$ well-ordered if either
$\xi_{1}\geq\xi_{2},\gamma_{1}\geq\gamma_{2},\zeta_{1}\geq\zeta_{2}\text{ or }\xi_{1}\leq\xi_{2},\gamma_{1}\leq\gamma_{2},\zeta_{1}\leq\zeta_{2}.$
In other words, two points are called well-ordered if all components of one of the points are not smaller than the corresponding coordinates of the other point.
If two points$(\xi_{1},\gamma_{1},\zeta_{1})$ and $(\xi_{2},\gamma_{2},\zeta_{2})$ are not well-ordered, then, without loss of generality, we can assume that:
$\xi_{1}\geq\xi_{2},\gamma_{1}>\gamma_{2},\zeta_{1}<\zeta_{2}.$
All other versions can be transferred into this, either by changing the names of variables or the names of points.
Now we present an important lemma, which is a direct consequence of (2.6).
Lemma 2.1.
For any function $f\in C(\Omega)$ satisfying (2.6) and any two points $x_{1}=(\xi_{1},\gamma_{1},\zeta_{1}),x_{2}=(\xi_{2},\gamma_{2},\zeta_{2})\in\Omega$ the inequality
$f(\xi_{1},\gamma_{1},\zeta_{1})+f(\xi_{2},\gamma_{2},\zeta_{2})\leq f\left((\xi_{1},\xi_{2}),(\gamma_{1},\gamma_{2})(\zeta_{1},\zeta_{2})\right)+f\left([\xi_{1},\xi_{2}],[\gamma_{1},\gamma_{2}],[\zeta_{1},\zeta_{2})]\right)$ (2.8)
holds, where we denote $(a,b)=\min(a,b)$ and $[a,b]=\max(a,b)$. The equality happens iff $x_{1}$ and $x_{2}$ are well-ordered.
Proof.
If $x_{1}$ and $x_{2}$ are well-ordered, then the expressions on the left and right-hand sides become the same, and the equality holds.
If $x_{1}$ and $x_{2}$ are not well-ordered, without loss of generality we can assume that $\xi_{1}\geq\xi_{2}$, $\gamma_{1}>\gamma_{2}$ and $\zeta_{1}<\zeta_{2}$. Thus, we have
$(\xi_{1},\xi_{2})=\xi_{2},(\gamma_{1},\gamma_{2})=\gamma_{2},(\zeta_{1},\zeta_{2})=\zeta_{1}\text{ and }[\xi_{1},\xi_{2}]=\xi_{1},[\gamma_{1},\gamma_{2}]=\gamma_{1},[\zeta_{1},\zeta_{2}]=\zeta_{2}.$
Let’s name $x_{1}^{\prime}=(\xi_{1},\gamma_{1},\zeta_{2}),x_{2}^{\prime}=(\xi_{2},\gamma_{2},\zeta_{1})$, $x_{1}^{\prime\prime}=(\xi_{1},\gamma_{2},\zeta_{1})$ and $x_{2}^{\prime\prime}=(\xi_{1},\gamma_{2},\zeta_{2})$. The inequality is equivalent to
$f(x_{1}^{\prime})+f(x_{2}^{\prime})-f(x_{1})-f(x_{2})$
$=\left(f(x_{2}^{\prime})+f(x_{2}^{\prime\prime})-f(x_{1}^{\prime\prime})-f(x_{2})\right)+\left(\left(f(x_{1}^{\prime\prime})+f(x_{1}^{\prime})-f(x_{2}^{\prime\prime})-f(x_{1})\right)\right)>0$
Which, directly follows from (2.6)(See: Figure 1). ∎
$x$$z$$y$$x_{2}^{\prime\prime}$$x_{1}^{\prime\prime}$$x_{2}^{\prime}$$x_{2}$$x_{1}$$x_{1}^{\prime}$
Figure 1: Geometric interpretation of the proof of Lemma 2.1
Definition 2.2.
A projection cycle-vector $(p^{0},\lambda^{0})\in X_{S}$ is called optimal for the function $f$ on the finite set $S$ if
$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
The set of all optimal projection cycle-vectors for $f$ on $S$ is denoted by $X_{S,f}$.
Obviously, for any $(p,\lambda)\in X_{S,f}$ we have
$\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq 0$ (2.9)
as otherwise, taking $(p,-\lambda)$ will contradict the optimality of $(p,\lambda)$.
Definition 2.3.
The sum of two projection cycles-vectors is a projection cycle-vector on the union of the points of the given projection cycle-vectors, with weight at each point equal to the sum of the weights of the given projection cycle-vectors on that point.
We can express the above definition by formulas as follows: $(p^{1},\lambda^{1})+(p^{2},\lambda^{2})=(p,\lambda)$ where, $p=p^{1}\cup p^{2}$ and
$\lambda(x)=\begin{cases}\lambda^{1}(x),\hskip 45.0ptx\in p^{1}\backslash p^{2}\\ \lambda^{1}(x)+\lambda^{2}(x),\hskip 8.0ptx\in p^{1}\cap p^{2}\\ \lambda^{2}(x),\hskip 43.0ptx\in p^{2}\backslash p^{1}\\ \end{cases}$ (2.10)
for any two (weak) projection cycle vectors $(p^{1},\lambda^{1})$ and $(p^{2},\lambda^{2})$.
Clearly,if both $(p^{1},\lambda^{1})\in X_{S}$ and $(p^{2},\lambda^{2})\in X_{S}$ for some finite set $S$ and function $f$, then
$(p^{1},\lambda^{1})+(p^{2},\lambda^{2})\in X_{S}$
since $p^{1}\subset S$ and $p^{2}\subset S$ is equivalent to $p^{1}\cup p^{2}\subset S$. Moreover, we have
$\sum_{x_{i}\in p}\lambda_{i}f(x_{i})=\sum_{x_{j}\in p^{1}}\lambda^{1}_{j}f(x_{j})+\sum_{x_{k}\in p^{2}}\lambda^{2}_{k}f(x_{k})$
Lemma 2.2.
Let $(p^{1},\lambda^{1})$, and $(p^{2},\lambda_{\varepsilon})$ be two projection cycle vectors, where, $\lambda_{\varepsilon}$ is a vector of length $m$ such that each component is $\varepsilon$ or $-\varepsilon$ for some positive real number $\varepsilon$. Let $k$ be the number of common points of $p^{1}$ and $p^{2}$, such that the corresponding components of the vectors $\lambda$ and $\lambda_{\varepsilon}$ have opposite signs. If $k\geq\frac{m}{2}$, then for
$0<\varepsilon\leq\min_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda^{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}|\lambda^{1}(x)|$
we have
$\sum_{x\in p}|\lambda(x)|\leq\sum_{x\in p^{1}}|\lambda^{1}(x)|$ (2.11)
where, $(p,\lambda)=(p^{1},\lambda^{1})+(p^{2},\lambda_{\varepsilon})$.
Proof.
By the definition (2.10) we have,
$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|=\sum_{x\in p^{1}\cap p^{2}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon$
The first summand can be written as,
$\sum_{x\in p^{1}\cap p^{2}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)=$
$=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)$
By the triangle inequality, we have
$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)\leq\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}|\lambda_{\varepsilon}(x)|=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon$
On the other hand, for
$0<\varepsilon\leq\min_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}|\lambda^{1}(x)|$
we have
$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)=-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon$
Thus, we have
$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|\leq-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon\leq 0$
Where the last inequality follows from the fact that the number of summands with a negative sign is $k$, the number of all summands is $m$ and $k\geq\frac{m}{2}$
The next theorem deals with the geometric structure of optimal projection cycles.
Theorem 2.2.
Let $(p,\lambda)$ be a projection cycle-vector, that is optimal for a function $f\in C(\Omega)$ satisfying (2.6) on the set $S=S_{x}\times S_{y}\times S_{z}$ for some finite sets $S_{x},S_{y},S_{z}\subset[0,1]$ with $\{0,1\}\subset S_{x}\cap S_{y}\cap S_{z}$. Then the following propositions hold:
  1. (a)
    If $\lambda_{k},\lambda_{j}$ are both positive for some indices $k,j$ then, $x_{k}$ and $x_{j}$ are well-ordered,
  2. (b)
    Let’s call a plane on $\mathbb{R}^{3}$ an interior plane of the unit cube if it is parallel to one of the faces of the unit cube and passes through an interior point of the cube. If $p$ contains a well-ordered pair of points $x_{i},x_{j}$ lying on the same interior plane, then, $\lambda_{i},\lambda_{j}>0$,
  3. (c)
    If $\lambda_{i}<0$, then the point $x_{i}$ belongs to one of the edges of the unit cube not containing any of $(0,0,0)$ and $(1,1,1)$,
  4. (d)
    If any point $x_{i}$ of $p$, other than $(0,0,0)$ and $(1,1,1)$, belongs to one of the edges of the unit cube, then $\lambda_{i}<0$,
  5. (e)
    $(0,0,0),(1,1,1)\in p$ and, both has positive weights.
Proof.
(a) Assume the contrary: Let, $(p,\lambda)=((x_{0},x_{1},...,x_{m}),(\lambda_{0},\lambda_{1},...,\lambda_{m}))\in X_{S,f}$ contain a non well-ordered pair of points $x_{k},x_{j}$ with $\lambda_{k}>0,\lambda_{j}>0$.
Let the coordinates of the points $x_{k},x_{j}$ be $(\xi_{k},\gamma_{k},\zeta_{k})$ and $(\xi_{j},\gamma_{j},\zeta_{j})$ respectively. Since $x_{k},x_{j}$ are not well-ordered, without loss of generality we can assume that $\xi_{k}\geq\xi_{j}$, $\gamma_{k}>\gamma_{j}$ and $\zeta_{k}<\zeta_{j}$.
Denote the points $(\xi_{k},\gamma_{k},\zeta_{j})$ and $(\xi_{j},\gamma_{j},\zeta_{k})$ by $x_{k}^{\prime}$ and $x_{j}^{\prime}$ respectively. Clearly, $x_{k}^{\prime},x_{j}^{\prime}\in S$ and the projection cycle-vector $(p_{0},\lambda_{\varepsilon})\in X_{S}$ with $p_{0}=(x_{k},x_{j},x_{k}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,-\varepsilon,\varepsilon,\varepsilon)$ is well defined for any $\varepsilon>0$. Let, $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$.
Since $\lambda(x_{k})$ and $\lambda(x_{j})$ are positive, when $\varepsilon>0$ is small enough by Lemma $2.1$ we have $0\leq\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$.
On the other hand,
$\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})=\varepsilon(f(x^{\prime}_{k})+f(x^{\prime}_{j})-f(x_{k})-f(x_{j}))>0$
where the last inequality follows by Lemma 2.1 applied to the points $x_{k}$ and $x_{j}$ and the positivity of $\varepsilon$.
Now since, $\sum\lambda^{\prime}_{i}f(x_{i})>\sum\lambda_{i}f(x_{i})\geq 0$ and $0<\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$, we have
$\frac{\sum\lambda^{\prime}_{i}f(x_{i})}{\sum|\lambda^{\prime}_{i}|}-\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\frac{\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}>0$
Which contradicts to the optimality of $p,\lambda)$.
(b) Let $(p,\lambda)=((x_{0},x_{1},...,x_{m}),(\lambda_{0},\lambda_{1},...,\lambda_{m}))$ be an optimal projection cycle-vector for $f$ on the set $S$ that contains a well-ordered pair of points $x_{i},x_{j}$ from the same interior plane.
Without loss of generality we can assume that, $x_{i}$ and $x_{j}$ are lying on the same plane paralel to the $xOy$ plane and thus, we can write, $x_{i}=(\xi_{i},\gamma_{i},\zeta)$ and $x_{j}=(\xi_{j},\gamma_{j},\zeta)$ for some numbers $\xi_{i},\xi_{j},\gamma_{i},\gamma_{j}\in[0,1]$ and some $\zeta\in(0,1)$.
Assume $\xi_{i}\leq\xi_{j}$, then, since $x_{i},x_{j}$ are well-ordered we have $\gamma_{i}\leq\gamma_{j}$. If both $\lambda_{i},\lambda_{j}$ are not positive, have three cases:
  1. Case 1:
    $\lambda_{i}>0>\lambda_{j}$
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},0)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},0)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$
  2. Case 2:
    $\lambda_{i}<0<\lambda_{j}$
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},1)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},1)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,-\varepsilon,-\varepsilon,\varepsilon)$
  3. Case 3:
    $\lambda_{i},\lambda_{j}<0$
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{j},\zeta)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{i},\zeta)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$
$x$$z$$y$$x_{i}^{\prime}$$x_{i}$$x_{j}^{\prime}$$x_{j}$
Figure 2: Case 1: We simply replace,the pair $x_{i},x_{j}$ with $x^{\prime}_{i},x^{\prime}_{j}$
In each case we have chosen $\lambda_{\varepsilon}$ in a way that, at both of the points $x_{i}$ and $x_{j}$ the projection cycle-vectors $(p,\lambda)$ and $(p_{0},\lambda_{\varepsilon})$ have weights with opposite signs. Thus, by the Lemma $2.2$, for $\varepsilon>0$ small enough, we have
$0<\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|,$
where $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$.
On the other hand, applying the Lemma 2.1 gives
$\sum_{x\in p^{\prime}}\lambda^{\prime}(x)f(x)\geq\sum_{x\in p}\lambda(x)f(x).$
Hence, in all three cases, we will get a contradiction to the optimality of $(p,\lambda)$ simultaneously as in part $(a)$.
(c) Let $(p,\lambda)=(x_{0},x_{1},...,x_{m},\lambda_{0},\lambda_{1},...,\lambda_{m})$ be an optimal projection cycle-vector for $f$ on the set $S$ that contains a point $x_{i}=(\xi_{i},\gamma_{i},\zeta_{i})$ with $\lambda_{i}<0$.
Consider the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$. By the definition of the projection cycle each of these planes contains at least one point of $p$ with positive weight. Let these points be $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j}),x_{l}=(\xi_{l},\gamma_{i},\zeta_{l}),x_{k}=(\xi_{k},\gamma_{k},\zeta_{i})$. If any of these positive-weighted points makes a well-ordered pair with $x_{i}$, then the projection cycle-vector $(p,\lambda)$ cannot be optimal by part $(b)$.
$x$$z$$y$$x_{i}$
Figure 3: The possible locations of the points that balances $x_{i}$
due to part $(b)$.
Assume both $10<\xi<$ 1 and $10<\gamma<$ 1. By the part $(b)$ we have $(x_{i},x_{j})$ is not a well-ordered pair,thus, either $\gamma_{i}<\gamma_{j},\zeta_{i}>\zeta_{j}$ or $\gamma_{i}>\gamma_{j},\zeta_{i}<\zeta_{j}$. Similarly, $(x_{i},x_{l})$ is not a well-ordered pair, thus, either $\xi_{i}<\xi_{l},\zeta_{i}>\zeta_{l}$ or $\xi_{i}>\xi_{l},\zeta_{i}<\zeta_{l}$. It’s easy to verify that in each of the four possible cases, $x_{j},x_{l}$ cannot be a well-ordered pair. By part $(a)$ if $(p,\lambda)$ is optimal then any pair of points $x_{j},x_{l},x_{k}$ should be well-ordered. Thus, $(p,\lambda)$ cannot be optimal, and we get a contradiction. Since, we choose $\xi,\gamma$ arbitrarily from $\xi,\gamma,\zeta$, this contradiction works if at least two of the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$ are inner planes, which happens when $x_{i}$ is not on the edges of the unit cube. Hence, $x_{i}$ belongs to one of the edges of the unit cube.
Now, if, $x_{i}$ and $(0,0,0)$ are on the same edge, we have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$. Assume, $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist an other point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}>0$. Moreover, since, $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered. But this contradicts the optimality of $(p,\lambda)$, by part $(b)$. Similarly, in the other two cases, we get a contradiction.
Using similar arguments, it can be proven that $x_{i}$ and $(1,1,1)$ cannot be on the same edge.
Thus, $x_{i}$ belongs to one of the edges of the cube, not containing any of the vertices $(0,0,0)$ and $(1,1,1)$. Let’s name these edges $l_{1},...,l_{6}$ as following
$l_{1}=\{(\xi,0,1),0\leq\xi\leq 1\}$
$l_{2}=\{(1,0,\zeta),0\leq\zeta\leq 1\}$
$l_{3}=\{(1,\gamma,0),0\leq\gamma\leq 1\}$
$l_{4}=\{(\xi,1,0),0\leq\xi\leq 1\}$
$l_{5}=\{(0,1,\zeta),0\leq\zeta\leq 1\}$
$l_{6}=\{(0,\gamma,1),0\leq\gamma\leq 1\}$
(d) Assume the contrary, let $x_{i}\in p$ be a point on one of the edges of the unit cube with $\lambda_{i}>0$. We have three cases: Case 1: $x_{i}$ is on the same edge with one of the vertices $(0,0,0)$ or $(1,1,1)$.
If, $x_{i}$ and $(0,0,0)$ are on the same edge, have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$. Assume, $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist an other point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}<0$. Moreover, since, $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered. But this contradicts the optimality of $(p,\lambda)$, by part $(b)$.
Similarly, in the other two cases, we get a contradiction. Using similar arguments, it can be proven that $x_{i}$ and $(1,1,1)$ cannot be on the same edge.
Case 2: $x_{i}$ belongs to one of the edges of the cube, not containing any of the vertices $(0,0,0)$ and $(1,1,1)$. Without loss of generality, we can assume that $x_{i}=(\xi,0,1)$. To make the sum of weights on the plane $z=1$ equal to zero, there should exist a point $x_{j}$ of $p$ with a negative weight on $z=1$. By the part $(c)$ the point $x_{j}$ belongs to at least one of the edges $l_{1}$ or $l_{6}$. Thus, we have two sub-cases:
Case 2.1: $x_{j}=(\xi_{j},0,1)$ for some $0\leq\xi_{j}\leq 1$ Clearly, $\xi_{j}\neq\xi$. We should consider two sub-cases:
Case 2.1.1: $\xi_{j}<\xi$
Consider the projection cycle $(p,\lambda_{\varepsilon})$, where $p_{0}=(x_{i},x_{j},(\xi,1,1),(\xi_{j},1,1))$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$
Case 2.1.2: $\xi<\xi_{j}$
Consider the projection cycle $(p_{0},\lambda_{\varepsilon})$, where $p_{0}=(x_{i},x_{j},(\xi,0,0),(\xi_{j},0,0))$ and $\lambda_{\varepsilon}=(\varepsilon,-\varepsilon,-\varepsilon,\varepsilon)$.
By a similar way to the part $(a)$ and part $(b)$, taking $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$ we get a contradiction to the optimality of $(p,\lambda)$.
Case 2.2: $x_{j}=(0,\gamma,1)$ for some $0<\gamma\leq 1$ Consider the plane, $y=\gamma$. To make the sum of weights on the plane $y=\gamma$ equal to zero, there should exist a point $x_{k}$ of $p$ with a positive weight on $y=\gamma$. Let $x_{k}=(\xi_{k},\gamma,\zeta)$.If $\zeta=1$, the points, $x_{j}$ and $x_{k}$ belonging to the inner plane $y=\gamma$ becomes well-ordered and since $x_{j}$ and $x_{k}$ have opposite weights we get a contradiction by part $(b)$. If, $\zeta<1$, the points, $x_{i}$ and $x_{k}$ becomes non well-ordered. Since both of $x_{i}$ and $x_{k}$ have positive weights, we get a contradiction by the part $(a)$.
Since, we get a contradiction in each case, the proof of part $(d)$ is completed.
(e) Consider the point $x_{i}=(\xi_{i},\gamma_{i},\zeta_{i})\in p$ with the minimal coordinates among positive weighted points of $p$. (Since, the positive weighted points are well-ordered pairwise, the point with the minimal sum of coordinates has minimal coordinates.)
If $x_{i}=(0,0,0)$, we are done. Otherwise, since $x_{i}$ is chosen to be minimal, it cannot be any other vertex of the unit cube, and thus there is at least one inner plane passing through $x_{i}$. Without loss of generality, assume this plane has equation $x=\xi_{i}$. Then, on the plane $x=\xi_{i}$ there is another point $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j})\in p$ such that $\lambda_{j}<0$. By part $(b)$ we know that $x_{i}$ and $x_{j}$ cannot be well-ordered and hence, either $\gamma_{j}<\gamma_{i}$ or $\zeta_{j}<\zeta_{i}$ but in the first case on the plane $y=\gamma_{j}$ and in the second case on the plane $z=\zeta_{j}$ there is a positive weighted point to make the sum over that plane equal to zero. But this contradicts to the minimality of $x_{i}$ and hence, $x_{i}=(0,0,0)$.
Similarly, we can prove that the positive weighted point with maximal coordinates is $(1,1,1)$.
$x$$z$$y$
Figure 4: Geometric interpretation: Theorem 2.2
Let $X^{\prime}_{S}$ be the set of projection cycle-vectors $(p,\lambda)\in X_{S}$, such that the propositions $(a)-(e)$ of the Theorem 2.2 holds for $(p,\lambda)$ and let $X^{\prime}_{S,min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in X^{\prime}_{S}$.
The following lemmas hold.
Lemma 2.3.
The vector assigned to a minimal projection cycle is unique up to multiplication by a constant. Moreover, if it is orthonormal, then it has rational components. Thus, any minimal projection cycle uniquely (up to a sign) defines an orthonormal projection cycle-vector.
Proof.
(See: Lemma 4 and Lemma 5 in [19] or Lemma 3.32 in [11]). ∎
Lemma 2.4.
For any projection cycle $(p,\lambda)\in X_{S}$ and for any function $f\in C(\Omega)$, there exist a projection cycle $(p^{0},\lambda^{0})\in X_{S,min}$ such that $p^{0}\subset p$ and the following relations hold
  1. 1.
    $\lambda^{0}(x)\lambda(x)>0,\forall x\in p^{0}$
  2. 2.
    $\frac{\sum_{x_{j}\in p^{0}}\lambda^{0}_{j}f(x_{j})}{\sum|\lambda^{0}_{j}|}\geq\frac{\sum_{x_{i}\in p}\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
Proof.
See [19, Theorem 2]. ∎
The Lemma 2.3 implies that, for any function $f$,
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
and
$\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
Lemma 2.5.
For any set $S=S_{x}\times S_{y}\times S_{z}$ where, $S_{x},S_{y},S_{z}\subset[0,1]$ are some finite sets with $\{0,1\}\subset S_{x}\cap S_{y}\cap S_{z}$ and for any function $f\in C(\Omega)$ satisfying the inequalities (2.1), the intersection $X^{\prime}_{S}\cap X_{S,f}$ is non-empty.
Proof.
Let’s define
$f_{n}(x,y,z)=f(x,y,z)+\frac{1}{n}(xy+xz+yz),\forall n\in\mathbb{N}$
Clearly, the function $f_{n}$ satisfies the inequalities (2.6) and therefore it follows from Theorem 2.2 and Lemma 2.4 that,
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}$
Here, we can change the order of the limit and the maximum as $X_{S,min}$ is a finite set by the Lemma 2.3 (considering vectors up to multiplication by a constant) and hence,
$\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}$
By the Theorem 2.2 and Lemma 2.4,
$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}$
Changing the order of the limit and maximum once more and using the Lemma 2.3 gives the desired result:
$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}$
Now let’s choose $(p^{0},\lambda^{0})\in X^{\prime}_{S}$ such that,
$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}.$
Then,
$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}$
and thus, $(p^{0},\lambda^{0})\in X_{S,f}$ by the definition of $X_{S,f}$. Hence, $(p^{0},\lambda^{0})\in X^{\prime}_{S}\cap X_{S,f}$
2.2 Construction of a finite set containing an extremal projection cycle-vector
Above we study some structural properties of optimal projection cycles on certain finite sets. Below, we will use this properties to prove that there is a projection cycle on an explicitly given finite set that gives the supremum (2.4).
We start by defining the desired set. Let’s name the corners of the unit cube from $T_{1}$ to $T_{8}$ as
$T_{1}=(0,0,0),T_{2}=(0,0,1),T_{3}=(1,0,1),T_{4}=(1,0,0),$
$T_{5}=(1,1,0),T_{6}=(0,1,0),T_{7}=(0,1,1),T_{8}=(1,1,1).$
Define the functions,
$g_{1}(x,y,z)=f(x,y,z)-f(x,0,1)-f(1,y,0)-f(0,1,z),$
$g_{2}(x,y,z)=f(x,y,z)-f(1,0,z)-f(x,1,0)-f(0,y,1),$
$g_{3}(x,y,z)=f(x,y,z)-f(x,0,1)-f(1,0,z)-f(1,y,0),$
$g_{4}(x,y,z)=f(x,y,z)-f(1,0,z)-f(1,y,0)-f(x,1,0),$
$g_{5}(x,y,z)=f(x,y,z)-f(1,y,0)-f(x,1,0)-f(0,1,z),$
$g_{6}(x,y,z)=f(x,y,z)-f(x,1,0)-f(0,1,z)-f(0,y,1),$
$g_{7}(x,y,z)=f(x,y,z)-f(0,1,z)-f(0,y,1)-f(x,0,1),$
$g_{8}(x,y,z)=f(x,y,z)-f(0,y,1)-f(x,0,1)-f(1,0,z),$
on the unit cube and
$h_{1}(x,y,0)=f(x,y,0)-f(0,y,1)-f(x,0,1),$
$h_{2}(x,y,1)=f(x,y,1)-f(1,y,0)-f(x,1,0),$
$h_{3}(x,0,z)=f(x,0,z)-f(0,1,z)-f(x,1,0),$
$h_{4}(x,1,z)=f(x,1,z)-f(1,0,z)-f(x,0,1),$
$h_{5}(0,y,z)=f(0,y,z)-f(1,0,z)-f(1,y,0),$
$h_{6}(1,y,z)=f(1,y,z)-f(0,1,z)-f(0,y,1).$
on the faces of the unit cube lying on the planes $z=0,z=1,y=0,y=1,x=0$ and $x=1$ correspondingly.
Let $M_{i},1\leq i\leq 8$ be a maximum point of the function $g_{i}$ on the domain $\Omega=[0,1]^{3}$. And $F_{i},1\leq i\leq 6$ be a maximum point of the function $h_{i}$ on the face it’s defined.
Let
$U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$
and let $U_{x},U_{y},U_{z}$ be the set of projections of elements of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively. We will see that, the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ will not change the correctness of the following theorem.
Remark 2.2.
[…]
Theorem 2.3.
Let $f\in C(\Omega)$ be a function satisfying the inequalities (2.1) and $S=S_{x}\times S_{y}\times S_{z}\subset\Omega$ be any arbitrary finite set with $U\subset S$, then we have:
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$ (2.12)
Proof.
Let, $X^{\prime\prime}_{S}$ be the set of projection cycle-vectors $(p,\lambda)\in X_{S}$ that satisfies the propositions $(b),(c),(d)$ of the Theorem 2.2.
It is clear that, $X^{\prime}_{S}\subset X^{\prime\prime}_{S}$ and hence from the Lemma 2.5 we have,
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
Thus, it is enough to prove that
$\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
Assume the contrary. Since, $U\subset S$ and $S=S_{x}\times S_{y}\times S_{z}$ we have $U_{x}\subset S_{x}$, $U_{y}\subset S_{y}$ and $U_{z}\subset S_{z}$ which imply $U_{x}\times U_{y}\times U_{z}\subset S$ and therefore we can write the inequality
$\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
Hence, there exist a projection cycle $(p^{0},\lambda^{0})\in X^{\prime\prime}_{S}$ such that
$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.13)
Clearly, $(p^{0},\lambda^{0})\not\in X_{U_{x}\times U_{y}\times U_{z}}$.
If all points in $p^{0}$ with positive weight belong to $U$, then the set of the projections of them into the edges of the unit cube (which are all the possible locations of points of $p^{0}$ with negative weights) would belong to the set $U_{x}\times U_{y}\times U_{z}$ and hence we would have $p^{0}\subset U_{x}\times U_{y}\times U_{z}$.
Thus, there exist a point $x\in p^{0}$ with positive weight($\lambda^{0}(x)>0$) such that, $x\not\in U$. Moreover, let $(p^{0},\lambda^{0})$ be chosen to minimise the number of positive weighted points not belonging to the set $U$ among all projection cycle-vectors $(p,\lambda)\in X^{\prime\prime}_{S}$ satisfying (2.13).
By and Lemma 2.5 and Theorem 2.2 part $(d)$, we have that $x$ cannot be on the edges of the unit cube. Thus, we have the following cases:
$x$ is an interior point of one of the faces of the unit cube $\Omega$.
Without loss of generality, we can assume that $x$ belongs to the face lying on the $xOy$ plane. Then we can write $x=(\xi,\gamma,0)$ for some $\xi,\gamma\in(0,1)$.
Consider the planes $x=\xi$ and $y=\gamma$. In each of these planes, at least one point of $(p^{0},\lambda^{0})$ with negative weight should exist. By Theorem 2.2, part $(c)$, these negative weighted points could only be on the edges, not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$. The intersections of the planes $x=\xi$ and $y=\gamma$ and the edges are the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ respectively.
Since both of $(\xi,1,0)$ and $(1,\gamma,0)$ are well-ordered with $x$ and lying on an interior plane, by the Theorem 2.2, part $(b)$, none of $(\xi,1,0)$ and $(1,\gamma,0)$ can have a negative weight. Thus, both $(\xi,0,1)$ and $(0,\gamma,1)$ have negative weights. Moreover, since no other negative weighted points can exist on the planes $x=\xi$ and $y=\gamma$, we have:
$|\lambda^{0}(\xi,0,1)|\geq\lambda^{0}(x)\text{ and }|\lambda^{0}(0,\gamma,1)|\geq\lambda^{0}(x)$
and thus, $\min(|\lambda^{0}(\xi,0,1)|,|\lambda^{0}(0,\gamma,1)|,|\lambda^{0}(x)|)=\lambda^{0}(x)$.
Let’s move the weights on the points $x$,$(\xi,0,1)$ and $(0,\gamma,1)$ to the points $F_{1}=(\xi_{1},\gamma_{1},0)$ and its two corresponding projections, namely $(\xi_{1},0,1)$ and $(0,\gamma_{1},1)$. Or in other words, we define
$p^{\prime}=(x,(\xi,0,1)\,(0,\gamma,1),(\xi_{1},\gamma_{1},0),(\xi_{1},0,1),(0,\gamma_{1},1))$
and vector
$\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$
where, $\varepsilon=\lambda^{0}(x)$. Let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon})$. Since, at each of the first three points of $p^{\prime}$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs and
$\varepsilon=\min(|\lambda^{0}(\xi,0,1)|,|\lambda^{0}(0,\gamma,1)|,|\lambda^{0}(x))|,$
by the Lemma 2.2 we have
$\sum|\lambda^{1}_{i}|\leq\sum|\lambda^{0}_{i}|$
On the other hand,
$\sum\lambda^{1}_{i}f(x_{i})-\sum\lambda^{0}_{i}f(x_{i})=$
$=f(\xi_{1},\gamma_{1},0)-f(\xi_{1},1,0)-f(1,\gamma_{1},0)-f(\xi,\gamma,0)+f(\xi,1,0)+f(1,\gamma,0)=h(\xi_{1},\gamma_{1},0)-h(\xi,\gamma,0)\geq 0$
where the inequality follows from the choice of $F_{1}=(\xi_{1},\gamma_{1},0)$ as maximum of the function $h_{1}$. Thus, we get
$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
Clearly, $(p^{1},\lambda^{1})\in X^{\prime\prime}_{S}$ as no negative weighted point became positive weighted and vice versa. Moreover, $\lambda^{1}(x)=\lambda^{0}(x)-\lambda_{\varepsilon}(x)=0$ and hence, $(p^{1},\lambda^{1})$ has less number of positive weighted points not belonging $U$ than $(p^{0},\lambda^{0})$ has. Thus, we get a contradiction to the choice of $(p_{0},\lambda_{0})$.
Case 2: $x$ is an interior point of the unit cube $\Omega$.
Let $x=(\xi,\gamma,\zeta)$. Consider the planes $x=\xi$, $y=\gamma$, $z=\zeta$. Their intersections with the edges of the unit cube that are not adjacent to the vertices not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$ will be the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ and $(1,0,\zeta),(0,1,\zeta)$ respectively. To make the sum over the corresponding plane zero, at least one (maybe both) of two points in each plane should belong to $(p^{0},\lambda^{0})$ with a negative weight.
Hence, all points of at least one of the eight sets
$G_{1}(x)=\{(\xi,0,1),(1,\gamma,0),(0,1,\zeta)\},$
$G_{2}(x)=\{(1,0,\zeta),(\xi,1,0),(0,\gamma,1)\},$
$G_{3}(x)=\{(\xi,0,1),(1,0,\zeta),(1,\gamma,0)\},$
$G_{4}(x)=\{(1,0,\zeta),(1,\gamma,0),(\xi,1,0)\},$
$G_{5}(x)=\{(1,\gamma,0),(\xi,1,0),(0,1,\zeta)\},$
$G_{6}(x)=\{(\xi,1,0),(0,1,\zeta),(0,\gamma,1)\},$
$G_{7}(x)=\{(0,1,\zeta),(0,\gamma,1),(\xi,0,1)\},$
$G_{8}(x)=\{(0,\gamma,1),(\xi,0,1),(1,0,\zeta)\}$
belong to $(p^{0},\lambda^{0})$ and have negative weights.
For each $G_{i}$ we proceed by moving the weights at point $x$ and elements of $G_{i}$ to $M_{i}$ and its corresponding projections on the edges. For example, if all elements of $G_{1}$ belong to $p^{0}$, then we construct $p^{\prime}=(x,(\xi,0,1),(1,\gamma,0),(0,1,\zeta),(\xi_{1},\gamma_{1},\zeta_{1}),(\xi_{1},0,1),(1,\gamma_{!},0),(0,1,\zeta_{1}))$ and vector $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon,-\varepsilon)$ where $M_{1}=(\xi_{1},\gamma_{1},\zeta_{1})$ is the maximum point of $g_{1}$ defined above and $\varepsilon>0$.
As in Case 1, let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon}).$ Since, at each of the first four points of $(p^{\prime},\lambda_{\varepsilon})$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs, by the Lemma 2.2 for $\varepsilon=\min(\lambda)$ we have
$\sum|\lambda_{i}^{1}|\leq\sum|\lambda_{j}^{0}|.$
Moreover,
$\sum\lambda^{1}_{i}f(x^{1}_{i})-\lambda^{0}_{i}f(x^{0}_{i})=g_{1}(M_{1})-g_{1}(x)\geq 0.$
Thus, we have
$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}$
As in the first case the last inequality contradicts our assumption of $(p^{0},\lambda^{0})$. Similarly for all other sets we will get a contradiction.
Since all cases lead to the contradiction, the theorem is proved.
The proof implies that, the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ does not affect to the correctness of the theorem.
Now we are ready to prove the Theorem $2.1$.
Proof (Theorem 2.1).
Applying the Theorem 2.3 in the relation (2.5) we get
$E(f,\Omega)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{(U_{x}\cup A)\times(U_{y}\cup B)\times(U_{z}\cup C)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=$
$\sup_{A,B,C}\left(\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\right)=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
Remark 2.3.
The proofs of the Theorem 2.3 and the Theorem 2.1 shows that, there exists a projection cycle-vector $(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}$ such that $(p,\lambda)$ satisfies all the propositions of the Theorem 2.2 and
$E(f,\Omega)=\frac{\sum_{x_{i}\in p}\lambda(x_{i})f(x_{i})}{\sum_{x_{i}\in p}|\lambda(x_{i})|}.$
Moreover, it can be reconstructed by only knowing the weights in the vertices and all positive weighted points.
3 Minimal Projection Cycles on a Finite Set
The relation (2.2) implies that, the error can be found by calculating a maximum over a finite set of the form $V=U_{x}\times U_{y}\times U_{z}$. In this section we will search for $X^{\prime}_{V,min}$.
Let $p=(x_{1},x_{2},...,x_{k})$ be an ordered finite subset of $\Omega$ with $k$ points. In order to find all vectors $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$, such that $(p,\lambda)$ is a projection cycle-vector, we have to solve the system of linear equations (1.3). Let this system be represented as
$A\lambda=0$ (3.1)
for a matrix $A$. Below, we will prove some necessary conditions for an ordered subset $p^{\prime}$ of $p$ to be an optimal projection cycle.
We start with a well-known algebraic definition and a lemma (See: Definition 1 and Proposition 7 in [22]).
Definition 3.1.
Let for any vector $v=(v_{1},...,v_{n})\in\mathbb{R}^{n}$ we define,
$\operatorname{supp}(v):=\left\{i\leq m:v_{i}\neq 0\right\}$
and call it the support of $v$. A solution $v_{0}$ of the homogeneous linear equation $Mv=0$ is called a minimal solution if the set $\operatorname{supp}(v_{0})$ is minimal, that is, there exist no other solution $v$ of the equation satisfying
$\operatorname{supp}(v)\varsubsetneqq\operatorname{supp}(v_{0})$
Lemma 3.1.
Let $M$ be a matrix then a solution $v_{0}$ of the equation $Mv=0$ which $i_{1},i_{2},...,i_{t}$-th components are nonzero and all other components are zero, is minimal, then the column system consisting $i_{1},i_{2},...,i_{k}$-th columns of the matrix $M$ is a minimal linearly dependent column system.
Lemma 3.2.
Let $p^{\prime}=(x_{i_{1}},x_{i_{2}},\dots x_{i_{t}})\subset p$ be a minimal projection cycle and $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$. Then the vector, $\lambda=(\lambda_{1},...,\lambda_{k})$ is a minimal solution of the equation (3.1), where,
$\lambda_{i}=\begin{cases}\lambda^{\prime}_{i},\hskip 13.0pt\text{ if }i\in\{i_{1},i_{2},\dots,i_{t}\}\\ 0,\hskip 17.0pt\text{ otherwise }\end{cases}$ (3.2)
Proof.
Assume the contrary, suppose $\lambda$ is not a minimal solution of (3.1). Then, there exist a solution $\lambda^{0}$ of (3.1) such that,
$\operatorname{supp}(\lambda^{0})\varsubsetneqq\operatorname{supp}(\lambda).$
Thus, there exist a nonempty set $\{j_{1},j_{2},...,j_{s}\}\varsubsetneqq\{1,2,...,t\}$ such that the $i_{j_{1}},i_{j_{2}},...,i_{j_{s}}$-th components of $\lambda^{0}$ are nonzero and all other components are zero. And, this implies that
$p^{\prime\prime}=(x_{i_{j_{1}}},x_{i_{j_{2}}},...,x_{i_{j_{s}}})\varsubsetneqq p^{\prime}$
is a projection cycle with vector $\lambda^{\prime\prime}=(\lambda^{0}_{i_{j_{1}}},\lambda^{0}_{i_{j_{2}}},...,\lambda^{0}_{i_{j_{s}}})$, which contradicts the minimality of $p^{\prime}$. ∎
The next theorem gives a necessary condition on minimality of a projection cycle.
Theorem 3.1.
Let, $a_{i}$ denote the $i$-th column of the matrix $A$ and $p^{\prime}=(x_{i_{1}},x_{i_{2}},\dots x_{i_{t}})\subset p$ be a minimal projection cycle, then the set of columns $\{a_{i_{1}},a_{i_{2}},\dots a_{i_{t}}\}$ is a minimal linearly dependent set of columns.
Proof.
Let $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$ and the vector $\lambda$ be defined as in (3.2). Then, by the Lemma 3.2, $\lambda$ is a minimal solution of (3.1). Thus, by the Lemma 3.1 the proof follows. ∎
We can construct a system of equation with variables being the weights at points of $U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$ and equations being equations on the faces $x=0$, $x=1$, $y=0$, $y=1$, $z=0$ and $z=1$ correspondingly, as all other variables in (3.1) can be defined uniquely by the Remark 2.2. It is easy to verify that, this new system will be
$Mv=0,$
where
[…]
and $v=(\lambda(T_{1}),\lambda(T_{8}),\lambda(M_{1}),\lambda(M_{2}),\dots,\lambda(M_{8}),\lambda(T_{2}),\lambda(T_{3}),\dots,\lambda(T_{7}),\lambda(F_{1}),\lambda(F_{2}),\dots,\lambda(F_{6}))^{\mathrm{T}}.$
Reducing the matrix $M$ to row echelon form by using the fact that the rows $r_{1},\dots,r_{6}$ of the matrix satisfies the identity $r_{1}+r_{2}=r_{3}+r_{4}=r_{5}+r_{6},$ gives $M\sim M^{\prime}$, where $M^{\prime}$ is the matrix given below
$\begin{pmatrix}1&0&-1&-1&-1&-2&-2&-1&0&0&1&1&1&0&0&0&1&0&-1&-2&-1&0\\ 0&1&-1&-1&0&-1&-2&-2&-1&0&0&1&1&1&0&0&0&1&0&-1&-2&-1\\ 0&0&0&0&1&2&1&-1&-2&-1&-1&-1&0&1&1&0&-2&0&2&2&0&-2\\ 0&0&0&0&-1&1&2&1&-1&-2&0&-1&-1&0&1&1&-2&-2&0&2&2&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\end{pmatrix}$
By the Lemma 2.5 we know that there exist an optimal projection cycle on $U_{x}\times U_{y}\times U_{z}$ that has $\lambda(T_{1})>0$, $\lambda(T_{8})>0$, $\lambda(M_{i})\geq 0$, $1\leq i\leq 8$, $\lambda(F_{i})\geq 0$, $1\leq i\leq 6$, $\lambda(T_{i})\leq 0$, $2\leq i\leq 7$ and $\lambda(F_{i})\lambda(F_{j})=0$ if $i\neq j$ and $j-i$ is even. Thus, its enough to find all minimal solutions of $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0.$
Assume $v$ is a minimal solution of the equation $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0$. The Theorem 3.1 implies that the system of columns of $M^{\prime}$ corresponding to the nonzero components of $v$ is a minimal linear dependent system. It is not hard to verify that there are $123$ minimal column systems containing the first two columns and satisfying the above conditions. And these are the ones given below. Note that we give only the weights on the corner points and the points with positive weights. The remaining points and their weights can be reconstructed as explained above.
1) Only $T_{i}$ terms
$\textbf{1.1}:T_{1},T_{8},T_{2},T_{5}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(T_{2})=-1,\quad\lambda(T_{5})=-1$
$\textbf{1.2}:T_{1},T_{8},T_{3},T_{6}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(T_{3})=-1,\quad\lambda(T_{6})=-1$
$\textbf{1.3}:T_{1},T_{8},T_{4},T_{7}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(T_{4})=-1,\quad\lambda(T_{7})=-1$
$\textbf{1.4}:T_{1},T_{8},T_{2},T_{4},T_{6}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=1,\quad\lambda(T_{2})=-1,\quad\lambda(T_{4})=-1,\quad\lambda(T_{6})=-1$
$\textbf{1.5}:T_{1},T_{8},T_{3},T_{5},T_{7}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=2,\quad\lambda(T_{3})=-1,\quad\lambda(T_{5})=-1,\quad\lambda(T_{7})=-1$
2) $T_{i}$ and one $M_{j}$
$\textbf{2.1}:T_{1},T_{8},M_{1}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(M_{1})=1$
$\textbf{2.2}:T_{1},T_{8},M_{2}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(M_{2})=1$
$\textbf{2.3}:T_{1},T_{8},M_{3},T_{2},T_{6}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(M_{3})=1,\quad\lambda(T_{2})=-1,\quad\lambda(T_{6})=-2$
$\textbf{2.4}:T_{1},T_{8},M_{3},T_{5},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{3})=1,\quad\lambda(T_{5})=-1,\quad\lambda(T_{7})=-2$
$\textbf{2.5}:T_{1},T_{8},M_{3},T_{6},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{3})=1,\quad\lambda(T_{6})=-1,\quad\lambda(T_{7})=-1$
$\textbf{2.6}:T_{1},T_{8},M_{4},T_{2},T_{6}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(M_{4})=1,\quad\lambda(T_{2})=-2,\quad\lambda(T_{6})=-1$
$\textbf{2.7}:T_{1},T_{8},M_{4},T_{2},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{4})=1,\quad\lambda(T_{2})=-1,\quad\lambda(T_{7})=-1$
$\textbf{2.8}:T_{1},T_{8},M_{4},T_{3},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{4})=1,\quad\lambda(T_{3})=-1,\quad\lambda(T_{7})=-2$
$\textbf{2.9}:T_{1},T_{8},M_{5},T_{2},T_{3}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{5})=1,\quad\lambda(T_{2})=-1,\quad\lambda(T_{3})=-1$
$\textbf{2.10}:T_{1},T_{8},M_{5},T_{2},T_{4}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(M_{5})=1,\quad\lambda(T_{2})=-2,\quad\lambda(T_{4})=-1$
$\textbf{2.11}:T_{1},T_{8},M_{5},T_{3},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{5})=1,\quad\lambda(T_{3})=-2,\quad\lambda(T_{7})=-1$
$\textbf{2.12}:T_{1},T_{8},M_{6},T_{2},T_{4}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(M_{6})=1,\quad\lambda(T_{2})=-1,\quad\lambda(T_{4})=-2$
$\textbf{2.13}:T_{1},T_{8},M_{6},T_{3},T_{4}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{6})=1,\quad\lambda(T_{3})=-1,\quad\lambda(T_{4})=-1$
$\textbf{2.14}:T_{1},T_{8},M_{6},T_{3},T_{5}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{6})=1,\quad\lambda(T_{3})=-2,\quad\lambda(T_{5})=-1$
$\textbf{2.15}:T_{1},T_{8},M_{7},T_{3},T_{5}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{7})=1,\quad\lambda(T_{3})=-1,\quad\lambda(T_{5})=-2$
$\textbf{2.16}:T_{1},T_{8},M_{7},T_{4},T_{5}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{7})=1,\quad\lambda(T_{4})=-1,\quad\lambda(T_{5})=-1$
$\textbf{2.17}:T_{1},T_{8},M_{7},T_{4},T_{6}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(M_{7})=1,\quad\lambda(T_{4})=-2,\quad\lambda(T_{6})=-1$
$\textbf{2.18}:T_{1},T_{8},M_{8},T_{4},T_{6}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(M_{8})=1,\quad\lambda(T_{4})=-1,\quad\lambda(T_{6})=-2$
$\textbf{2.19}:T_{1},T_{8},M_{8},T_{5},T_{6}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{8})=1,\quad\lambda(T_{5})=-1,\quad\lambda(T_{6})=-1$
$\textbf{2.20}:T_{1},T_{8},M_{8},T_{5},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{8})=1,\quad\lambda(T_{5})=-2,\quad\lambda(T_{7})=-1$
3) $T_{i}$ and two $M_{j}$
$\textbf{3.1}:T_{1},T_{8},M_{3},M_{6}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{3})=1,\quad\lambda(M_{6})=1$
$\textbf{3.2}:T_{1},T_{8},M_{4},M_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{4})=1,\quad\lambda(M_{7})=1$
$\textbf{3.3}:T_{1},T_{8},M_{5},M_{8}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{5})=1,\quad\lambda(M_{8})=1$
$\textbf{3.4}:T_{1},T_{8},M_{3},M_{4},T_{7}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{3})=1,\quad\lambda(M_{4})=1,\quad\lambda(T_{7})=-3$
$\textbf{3.5}:T_{1},T_{8},M_{3},M_{5},T_{2}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{3})=1,\quad\lambda(M_{5})=2,\quad\lambda(T_{2})=-3$
$\textbf{3.6}:T_{1},T_{8},M_{3},M_{5},T_{7}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{3})=2,\quad\lambda(M_{5})=1,\quad\lambda(T_{7})=-3$
$\textbf{3.7}:T_{1},T_{8},M_{3},M_{7},T_{5}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{3})=1,\quad\lambda(M_{7})=2,\quad\lambda(T_{5})=-3$
$\textbf{3.8}:T_{1},T_{8},M_{3},M_{7},T_{6}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{3})=2,\quad\lambda(M_{7})=1,\quad\lambda(T_{6})=-3$
$\textbf{3.9}:T_{1},T_{8},M_{3},M_{8},T_{6}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{3})=1,\quad\lambda(M_{8})=1,\quad\lambda(T_{6})=-3$
$\textbf{3.10}:T_{1},T_{8},M_{4},M_{5},T_{2}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{4})=1,\quad\lambda(M_{5})=1,\quad\lambda(T_{2})=-3$
$\textbf{3.11}:T_{1},T_{8},M_{4},M_{6},T_{2}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{4})=2,\quad\lambda(M_{6})=1,\quad\lambda(T_{2})=-3$
$\textbf{3.12}:T_{1},T_{8},M_{4},M_{6},T_{3}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{4})=1,\quad\lambda(M_{6})=2,\quad\lambda(T_{3})=-3$
$\textbf{3.13}:T_{1},T_{8},M_{4},M_{8},T_{6}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{4})=1,\quad\lambda(M_{8})=2,\quad\lambda(T_{6})=-3$
$\textbf{3.14}:T_{1},T_{8},M_{4},M_{8},T_{7}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{4})=2,\quad\lambda(M_{8})=1,\quad\lambda(T_{7})=-3$
$\textbf{3.15}:T_{1},T_{8},M_{5},M_{6},T_{3}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{5})=1,\quad\lambda(M_{6})=1,\quad\lambda(T_{3})=-3$
$\textbf{3.16}:T_{1},T_{8},M_{5},M_{7},T_{3}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{5})=2,\quad\lambda(M_{7})=1,\quad\lambda(T_{3})=-3$
$\textbf{3.17}:T_{1},T_{8},M_{5},M_{7},T_{4}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{5})=1,\quad\lambda(M_{7})=2,\quad\lambda(T_{4})=-3$
$\textbf{3.18}:T_{1},T_{8},M_{6},M_{7},T_{4}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{6})=1,\quad\lambda(M_{7})=1,\quad\lambda(T_{4})=-3$
$\textbf{3.19}:T_{1},T_{8},M_{6},M_{8},T_{4}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{6})=2,\quad\lambda(M_{8})=1,\quad\lambda(T_{4})=-3$
$\textbf{3.20}:T_{1},T_{8},M_{6},M_{8},T_{5}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{6})=1,\quad\lambda(M_{8})=2,\quad\lambda(T_{5})=-3$
$\textbf{3.21}:T_{1},T_{8},M_{7},M_{8},T_{5}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{7})=1,\quad\lambda(M_{8})=1,\quad\lambda(T_{5})=-3$
4) $T_{i}$ and three $M_{j}$
$\textbf{4.1}:T_{1},T_{8},M_{3},M_{5},M_{7}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{3})=1,\quad\lambda(M_{5})=1,\quad\lambda(M_{7})=1$
$\textbf{4.2}:T_{1},T_{8},M_{4},M_{6},M_{8}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{4})=1,\quad\lambda(M_{6})=1,\quad\lambda(M_{8})=1$
5) $T_{i}$ and one $F_{k}$
$\textbf{5.1}:T_{1},T_{8},T_{2},F_{2}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=1,\quad\lambda(T_{2})=-2,\quad\lambda(F_{2})=1$
$\textbf{5.2}:T_{1},T_{8},T_{3},F_{3}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=2,\quad\lambda(T_{3})=-2,\quad\lambda(F_{3})=1$
$\textbf{5.3}:T_{1},T_{8},T_{4},F_{6}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=1,\quad\lambda(T_{4})=-2,\quad\lambda(F_{6})=1$
$\textbf{5.4}:T_{1},T_{8},T_{5},F_{1}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=2,\quad\lambda(T_{5})=-2,\quad\lambda(F_{1})=1$
$\textbf{5.5}:T_{1},T_{8},T_{6},F_{4}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=1,\quad\lambda(T_{6})=-2,\quad\lambda(F_{4})=1$
$\textbf{5.6}:T_{1},T_{8},T_{7},F_{5}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=2,\quad\lambda(T_{7})=-2,\quad\lambda(F_{5})=1$
$\textbf{5.7}:T_{1},T_{8},T_{2},T_{4},F_{3}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(T_{2})=-2,\quad\lambda(T_{4})=-2,\quad\lambda(F_{3})=1$
$\textbf{5.8}:T_{1},T_{8},T_{2},T_{6},F_{5}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(T_{2})=-2,\quad\lambda(T_{6})=-2,\quad\lambda(F_{5})=1$
$\textbf{5.9}:T_{1},T_{8},T_{3},T_{5},F_{6}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(T_{3})=-2,\quad\lambda(T_{5})=-2,\quad\lambda(F_{6})=1$
$\textbf{5.10}:T_{1},T_{8},T_{3},T_{7},F_{2}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(T_{3})=-2,\quad\lambda(T_{7})=-2,\quad\lambda(F_{2})=1$
$\textbf{5.11}:T_{1},T_{8},T_{4},T_{6},F_{1}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=2,\quad\lambda(T_{4})=-2,\quad\lambda(T_{6})=-2,\quad\lambda(F_{1})=1$
$\textbf{5.12}:T_{1},T_{8},T_{5},T_{7},F_{4}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(T_{5})=-2,\quad\lambda(T_{7})=-2,\quad\lambda(F_{4})=1$
6) $T_{i}$, one $M_{j}$, one $F_{k}$
$\textbf{6.1}:T_{1},T_{8},M_{3},T_{2},F_{3}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{3})=1,\quad\lambda(T_{2})=-1,\quad\lambda(F_{3})=1$
$\textbf{6.2}:T_{1},T_{8},M_{3},T_{5},F_{6}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{3})=1,\quad\lambda(T_{5})=-1,\quad\lambda(F_{6})=1$
$\textbf{6.3}:T_{1},T_{8},M_{3},T_{6},F_{1}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{3})=2,\quad\lambda(T_{6})=-4,\quad\lambda(F_{1})=1$
$\textbf{6.4}:T_{1},T_{8},M_{3},T_{6},F_{6}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{3})=2,\quad\lambda(T_{6})=-2,\quad\lambda(F_{6})=1$
$\textbf{6.5}:T_{1},T_{8},M_{3},T_{7},F_{2}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{3})=2,\quad\lambda(T_{7})=-4,\quad\lambda(F_{2})=1$
$\textbf{6.6}:T_{1},T_{8},M_{3},T_{7},F_{3}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{3})=2,\quad\lambda(T_{7})=-2,\quad\lambda(F_{3})=1$
$\textbf{6.7}:T_{1},T_{8},M_{4},T_{2},F_{3}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{4})=2,\quad\lambda(T_{2})=-4,\quad\lambda(F_{3})=1$
$\textbf{6.8}:T_{1},T_{8},M_{4},T_{2},F_{6}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{4})=2,\quad\lambda(T_{2})=-2,\quad\lambda(F_{6})=1$
$\textbf{6.9}:T_{1},T_{8},M_{4},T_{3},F_{6}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{4})=1,\quad\lambda(T_{3})=-1,\quad\lambda(F_{6})=1$
$\textbf{6.10}:T_{1},T_{8},M_{4},T_{6},F_{1}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{4})=1,\quad\lambda(T_{6})=-1,\quad\lambda(F_{1})=1$
$\textbf{6.11}:T_{1},T_{8},M_{4},T_{7},F_{1}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{4})=2,\quad\lambda(T_{7})=-2,\quad\lambda(F_{1})=1$
$\textbf{6.12}:T_{1},T_{8},M_{4},T_{7},F_{4}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{4})=2,\quad\lambda(T_{7})=-4,\quad\lambda(F_{4})=1$
$\textbf{6.13}:T_{1},T_{8},M_{5},T_{2},F_{4}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{5})=2,\quad\lambda(T_{2})=-2,\quad\lambda(F_{4})=1$
$\textbf{6.14}:T_{1},T_{8},M_{5},T_{2},F_{5}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{5})=2,\quad\lambda(T_{2})=-4,\quad\lambda(F_{5})=1$
$\textbf{6.15}:T_{1},T_{8},M_{5},T_{3},F_{1}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{5})=2,\quad\lambda(T_{3})=-2,\quad\lambda(F_{1})=1$
$\textbf{6.16}:T_{1},T_{8},M_{5},T_{3},F_{6}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{5})=2,\quad\lambda(T_{3})=-4,\quad\lambda(F_{6})=1$
$\textbf{6.17}:T_{1},T_{8},M_{5},T_{4},F_{1}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{5})=1,\quad\lambda(T_{4})=-1,\quad\lambda(F_{1})=1$
$\textbf{6.18}:T_{1},T_{8},M_{5},T_{7},F_{4}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{5})=1,\quad\lambda(T_{7})=-1,\quad\lambda(F_{4})=1$
$\textbf{6.19}:T_{1},T_{8},M_{6},T_{2},F_{5}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{6})=1,\quad\lambda(T_{2})=-1,\quad\lambda(F_{5})=1$
$\textbf{6.20}:T_{1},T_{8},M_{6},T_{3},F_{2}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{6})=2,\quad\lambda(T_{3})=-4,\quad\lambda(F_{2})=1$
$\textbf{6.21}:T_{1},T_{8},M_{6},T_{3},F_{5}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{6})=2,\quad\lambda(T_{3})=-2,\quad\lambda(F_{5})=1$
$\textbf{6.22}:T_{1},T_{8},M_{6},T_{4},F_{1}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{6})=2,\quad\lambda(T_{4})=-4,\quad\lambda(F_{1})=1$
$\textbf{6.23}:T_{1},T_{8},M_{6},T_{4},F_{4}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{6})=2,\quad\lambda(T_{4})=-2,\quad\lambda(F_{4})=1$
$\textbf{6.24}:T_{1},T_{8},M_{6},T_{5},F_{4}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{6})=1,\quad\lambda(T_{5})=-1,\quad\lambda(F_{4})=1$
$\textbf{6.25}:T_{1},T_{8},M_{7},T_{3},F_{2}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{7})=1,\quad\lambda(T_{3})=-1,\quad\lambda(F_{2})=1$
$\textbf{6.26}:T_{1},T_{8},M_{7},T_{4},F_{2}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{7})=2,\quad\lambda(T_{4})=-2,\quad\lambda(F_{2})=1$
$\textbf{6.27}:T_{1},T_{8},M_{7},T_{4},F_{3}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{7})=2,\quad\lambda(T_{4})=-4,\quad\lambda(F_{3})=1$
$\textbf{6.28}:T_{1},T_{8},M_{7},T_{5},F_{4}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{7})=2,\quad\lambda(T_{5})=-4,\quad\lambda(F_{4})=1$
$\textbf{6.29}:T_{1},T_{8},M_{7},T_{5},F_{5}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{7})=2,\quad\lambda(T_{5})=-2,\quad\lambda(F_{5})=1$
$\textbf{6.30}:T_{1},T_{8},M_{7},T_{6},F_{5}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{7})=1,\quad\lambda(T_{6})=-1,\quad\lambda(F_{5})=1$
$\textbf{6.31}:T_{1},T_{8},M_{8},T_{4},F_{3}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{8})=1,\quad\lambda(T_{4})=-1,\quad\lambda(F_{3})=1$
$\textbf{6.32}:T_{1},T_{8},M_{8},T_{5},F_{3}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=4,\quad\lambda(M_{8})=2,\quad\lambda(T_{5})=-2,\quad\lambda(F_{3})=1$
$\textbf{6.33}:T_{1},T_{8},M_{8},T_{5},F_{6}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=5,\quad\lambda(M_{8})=2,\quad\lambda(T_{5})=-4,\quad\lambda(F_{6})=1$
$\textbf{6.34}:T_{1},T_{8},M_{8},T_{6},F_{2}:\quad\lambda(T_{1})=4,\quad\lambda(T_{8})=3,\quad\lambda(M_{8})=2,\quad\lambda(T_{6})=-2,\quad\lambda(F_{2})=1$
$\textbf{6.35}:T_{1},T_{8},M_{8},T_{6},F_{5}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=4,\quad\lambda(M_{8})=2,\quad\lambda(T_{6})=-4,\quad\lambda(F_{5})=1$
$\textbf{6.36}:T_{1},T_{8},M_{8},T_{7},F_{2}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=2,\quad\lambda(M_{8})=1,\quad\lambda(T_{7})=-1,\quad\lambda(F_{2})=1$
7) $T_{i}$, two $M_{j}$, one $F_{k}$
$\textbf{7.1}:T_{1},T_{8},M_{3},M_{4},F_{6}:\quad\lambda(T_{1})=6,\quad\lambda(T_{8})=5,\quad\lambda(M_{3})=2,\quad\lambda(M_{4})=2,\quad\lambda(F_{6})=3$
$\textbf{7.2}:T_{1},T_{8},M_{3},M_{5},F_{1}:\quad\lambda(T_{1})=7,\quad\lambda(T_{8})=8,\quad\lambda(M_{3})=2,\quad\lambda(M_{5})=4,\quad\lambda(F_{1})=3$
$\textbf{7.3}:T_{1},T_{8},M_{3},M_{5},F_{6}:\quad\lambda(T_{1})=8,\quad\lambda(T_{8})=7,\quad\lambda(M_{3})=4,\quad\lambda(M_{5})=2,\quad\lambda(F_{6})=3$
$\textbf{7.4}:T_{1},T_{8},M_{3},M_{7},F_{2}:\quad\lambda(T_{1})=8,\quad\lambda(T_{8})=7,\quad\lambda(M_{3})=2,\quad\lambda(M_{7})=4,\quad\lambda(F_{2})=3$
$\textbf{7.5}:T_{1},T_{8},M_{3},M_{7},F_{3}:\quad\lambda(T_{1})=7,\quad\lambda(T_{8})=8,\quad\lambda(M_{3})=4,\quad\lambda(M_{7})=2,\quad\lambda(F_{3})=3$
$\textbf{7.6}:T_{1},T_{8},M_{3},M_{8},F_{3}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=6,\quad\lambda(M_{3})=2,\quad\lambda(M_{8})=2,\quad\lambda(F_{3})=3$
$\textbf{7.7}:T_{1},T_{8},M_{4},M_{5},F_{1}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=6,\quad\lambda(M_{4})=2,\quad\lambda(M_{5})=2,\quad\lambda(F_{1})=3$
$\textbf{7.8}:T_{1},T_{8},M_{4},M_{6},F_{1}:\quad\lambda(T_{1})=7,\quad\lambda(T_{8})=8,\quad\lambda(M_{4})=4,\quad\lambda(M_{6})=2,\quad\lambda(F_{1})=3$
$\textbf{7.9}:T_{1},T_{8},M_{4},M_{6},F_{4}:\quad\lambda(T_{1})=8,\quad\lambda(T_{8})=7,\quad\lambda(M_{4})=2,\quad\lambda(M_{6})=4,\quad\lambda(F_{4})=3$
$\textbf{7.10}:T_{1},T_{8},M_{4},M_{8},F_{3}:\quad\lambda(T_{1})=7,\quad\lambda(T_{8})=8,\quad\lambda(M_{4})=2,\quad\lambda(M_{8})=4,\quad\lambda(F_{3})=3$
$\textbf{7.11}:T_{1},T_{8},M_{4},M_{8},F_{6}:\quad\lambda(T_{1})=8,\quad\lambda(T_{8})=7,\quad\lambda(M_{4})=4,\quad\lambda(M_{8})=2,\quad\lambda(F_{6})=3$
$\textbf{7.12}:T_{1},T_{8},M_{5},M_{6},F_{4}:\quad\lambda(T_{1})=6,\quad\lambda(T_{8})=5,\quad\lambda(M_{5})=2,\quad\lambda(M_{6})=2,\quad\lambda(F_{4})=3$
$\textbf{7.13}:T_{1},T_{8},M_{5},M_{7},F_{4}:\quad\lambda(T_{1})=8,\quad\lambda(T_{8})=7,\quad\lambda(M_{5})=4,\quad\lambda(M_{7})=2,\quad\lambda(F_{4})=3$
$\textbf{7.14}:T_{1},T_{8},M_{5},M_{7},F_{5}:\quad\lambda(T_{1})=7,\quad\lambda(T_{8})=8,\quad\lambda(M_{5})=2,\quad\lambda(M_{7})=4,\quad\lambda(F_{5})=3$
$\textbf{7.15}:T_{1},T_{8},M_{6},M_{7},F_{5}:\quad\lambda(T_{1})=5,\quad\lambda(T_{8})=6,\quad\lambda(M_{6})=2,\quad\lambda(M_{7})=2,\quad\lambda(F_{5})=3$
$\textbf{7.16}:T_{1},T_{8},M_{6},M_{8},F_{2}:\quad\lambda(T_{1})=8,\quad\lambda(T_{8})=7,\quad\lambda(M_{6})=2,\quad\lambda(M_{8})=4,\quad\lambda(F_{2})=3$
$\textbf{7.17}:T_{1},T_{8},M_{6},M_{8},F_{5}:\quad\lambda(T_{1})=7,\quad\lambda(T_{8})=8,\quad\lambda(M_{6})=4,\quad\lambda(M_{8})=2,\quad\lambda(F_{5})=3$
$\textbf{7.18}:T_{1},T_{8},M_{7},M_{8},F_{2}:\quad\lambda(T_{1})=6,\quad\lambda(T_{8})=5,\quad\lambda(M_{7})=2,\quad\lambda(M_{8})=2,\quad\lambda(F_{2})=3$
8) $T_{i}$ and two $F_{k}$
$\textbf{8.1}:T_{1},T_{8},F_{1},F_{2}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(F_{1})=1,\quad\lambda(F_{2})=1$
$\textbf{8.2}:T_{1},T_{8},F_{3},F_{4}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(F_{3})=1,\quad\lambda(F_{4})=1$
$\textbf{8.3}:T_{1},T_{8},F_{5},F_{6}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=1,\quad\lambda(F_{5})=1,\quad\lambda(F_{6})=1$
9) $T_{i}$, one $M_{j}$, two $F_{k}$
$\textbf{9.1}:T_{1},T_{8},M_{3},F_{3},F_{6}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{3})=2,\quad\lambda(F_{3})=1,\quad\lambda(F_{6})=1$
$\textbf{9.2}:T_{1},T_{8},M_{4},F_{1},F_{6}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{4})=2,\quad\lambda(F_{1})=1,\quad\lambda(F_{6})=1$
$\textbf{9.3}:T_{1},T_{8},M_{5},F_{1},F_{4}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{5})=2,\quad\lambda(F_{1})=1,\quad\lambda(F_{4})=1$
$\textbf{9.4}:T_{1},T_{8},M_{6},F_{4},F_{5}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{6})=2,\quad\lambda(F_{4})=1,\quad\lambda(F_{5})=1$
$\textbf{9.5}:T_{1},T_{8},M_{7},F_{2},F_{5}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{7})=2,\quad\lambda(F_{2})=1,\quad\lambda(F_{5})=1$
$\textbf{9.6}:T_{1},T_{8},M_{8},F_{2},F_{3}:\quad\lambda(T_{1})=3,\quad\lambda(T_{8})=3,\quad\lambda(M_{8})=2,\quad\lambda(F_{2})=1,\quad\lambda(F_{3})=1$
Remark 3.1.
For the above minimal projection cycles $(p,\lambda)$, there exists a function $f$, satisfying conditions (2.1), such that $E(f;\Omega)=\frac{\sum\lambda_{i}f(x_{i})}{\sum\lambda_{i}}$. For example,
for the minimal projection cycle
1.1. $T_{1},T_{8},T_{2},T_{5}:\ \lambda(T_{1})=1,\ \lambda(T_{8})=1,\ \lambda(T_{2})=-1,\ \lambda(T_{5})=-1$
this is the function $f(x,y,z)=xz$:
$E(f,\Omega)=\frac{1}{4}=\frac{\sum_{i=1,8,2,5}\lambda(T_{i})f(T_{i})}{\sum_{i=1,8,2,5}|\lambda(T_{i})|};$
for the minimal projection cycle
1.5. $T_{1},T_{8},T_{3},T_{5},T_{7}:\quad\lambda(T_{1})=1,\quad\lambda(T_{8})=2,\quad\lambda(T_{3})=-1,\quad\lambda(T_{5})=-1,\quad\lambda(T_{7})=-1$
this is the function $f(x,y,z)=xyz$:
$E(f,\Omega)=\frac{1}{3}=\frac{\sum_{i=1,8,3,5,7}\lambda(T_{i})f(T_{i})}{\sum_{i=1,8,3,5,7}|\lambda(T_{i})|};$
for the minimal projection cycle
2.11. $T_{1},T_{8},M_{5},T_{3},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{5})=1,\quad\lambda(T_{3})=-2,\quad\lambda(T_{7})=-1$
[…]
$E(f,\Omega)=\frac{7}{48}=\frac{\sum_{i=1,8,5,7}\lambda(T_{i})f(T_{i})+\lambda(M_{5})f(M_{5})+\sum_{i=3,4,5}\lambda(M_{5,l_{i}})f(M_{5,l_{i}})}{\sum_{i=1,8,5,7}|\lambda(T_{i})|+|\lambda(M_{5})|+\sum_{i=3,4,5}|\lambda(M_{5,l_{i}})|},$
where $M_{5}=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, $M_{5,l_{i}}=\Pr_{l_{i}}M_{5}$, $i=3,4,5$, etc.

### v2

@@
 Consider the following approximation problem of a continuous function of three variables by the sum of three continuous functions of one variable:
-$E(f,\Omega)=\inf||f(x,y,z)-\phi(x)-\psi(y)-\omega(z)||_{\infty}=?$
+[…]
 where $f(x,y,z)$ is a given continuous function defined on $\Omega=[0,1]^{3}$ and the infimum runs over all triplets of continuous functions $\phi(x),\psi(y),\omega(z)$ defined on the unit interval $[0,1]$.
@@
 1Institute of Mathematics and Mechanics, Baku, Azerbaijan
-1Center for Mathematics and its Applications, Khazar University, Baku, Azerbaijan
 e-mail: 1aliyevrashid@mail.ru, 2 quliyevvuqar.213@gmail.com, 3 amilcbiyev23@gmail.com
+Keywords: Best approximation; coordinate functions; projection cycle; Golomb’s formula; minimal projection cycle.
+2020 Mathematics Subject Classification: 41A50; 41A30
 1 Introduction
 1.1 Problem statement in a general form and previous results
-Let $f:X\to\mathbb{R}$ be a continuous function in the $n$ variables, where $X=X_{1}\times X_{2}\times\dots X_{n}$ is the Cartesian product of the compact Hausdorff spaces $n$.
-Consider the approximation of $f$ by means of $g_{i},i=1,\dots,n$, where $g_{i}$ is a continuous function on $X_{i}$ for $i=1,\dots,n$.
+Let $f:X\to\mathbb{R}$ be a continuous function of $n$ variables, where $X=X_{1}\times X_{2}\times\dots X_{n}$ is the Cartesian product of $n$ compact Hausdorff spaces.
+Consider the approximation of $f$ by sums of $g_{i},i=1,\dots,n$, where $g_{i}$ is a continuous function on $X_{i}$ for $i=1,\dots,n$.
 In other words, the problem asks for
-$E(f,X)=\inf_{(g_{1},g_{2},...,g_{n})}||f(x_{1},x_{2},\dots,x_{n})-\sum_{i=1}^{n}g_{i}(x_{i})||_{C(X)}=?$ (1.1)
+[…]
 where the infimum runs over all $n$-tuples of functions $(g_{1},g_{2},...,g_{n})$ satisfying $g_{i}\in C(X_{i}),i=1,\dots,n$.
 In general, this problem is related to Hilbert’s 13th problem.
-Using A.S.
-Kronrod’s [16] results, A.N.
+Using A.S.Kronrod’s [27] results, A.N.
 Kolmogorov [25] showed that any function $f\in C([0,1]^{n})$ is representable as a superposition of continuous functions of three variables.
-Developing this method, V.I.
-Arnold [1], [2] proved that each continuous function of three variables can be represented by superposition of continuous functions of two variables.
+Developing this method, V.I.Arnold [2], [3] proved that each continuous function of three variables can be represented by a superposition of continuous functions of two variables.
 This is a negative answer to Hilbert’s conjecture.
-Further, A.N.
-Kolmogorov [15] proved the following remarkable, stronger theorem in a simple, elementary and elegant way:
-Kolmogorov’s Theorem. There exist increasing functions $\phi_{pq}\in C([0,1])$, $p=1,...,n$, $q=1,...,2n+1$, such that an arbitrary given function
+Further, A.N.Kolmogorov [26] proved the following remarkable, stronger theorem in a simple, elementary and elegant way:
+Kolmogorov’s Theorem. There exist increasing functions $\phi_{pq}\in C([0,1])$, p=1,,np=1,...,n, q=1,,2n+1q=1,...,2n+1, such that an arbitrary given function
 $f(x_{1},...,x_{n})\in C([0,1]^{n})$
@@
 $f(x_{1},...,x_{n})=\sum_{q=1}^{2n+1}g_{q}\left(\sum_{p=1}^{n}\phi_{pq}(x_{p})\right),$
-where $g_{q}\in C([0,1]$ depend on $f$.
-Thus, every continuous function on the $n$ variables can be represented by superposition of continuous functions of one variable and the simplest function of two variables, namely, the sum of variables.
+where $g_{q}\in C([0,1])$ depend on $f$.
+Thus, every continuous function on the $n$ variables can be represented by a superposition of continuous functions of one variable and the simplest function of two variables, namely, the sum of variables.
 In addition to the problems of finding a precise expression for a given function in terms of a combination of functions of fewer variables, it is natural to consider the best approximation of functions of several variables by combinations of functions of fewer variables.
-These problems arise in various areas of mathematics, for example, in solving integral equations (see [5]), in solving functional equations (see [4]), in saving information (see [6]), in scaling matrices (see [8], [9]), etc.
-In space $C([0,1]^{n})$ this problem was first studied by S.
+These problems arise in various areas of mathematics, for example, in solving integral equations (see [10]), functional equations (see [9]), in saving information (see [11]), in scaling matrices (see [14], [15]), in isogeometric analysis (see [8]), etc.
+In the space $C([0,1]^{n})$, this problem was first studied by S.
 P.
@@
 O’Farrell in [28].
-Diliberto and Straus gave a ’leveling algorithm’ to find the approximation error (1.1) for $n=2$ and propose a theorem that will ’analogously’ hold for $n\geq 3$.
+In [11], Diliberto and Straus gave a “leveling algorithm” to find the approximation error (1.1) for $n=2$.
+The addressed problem, together with the methods developed from its analysis, finds important applications in approximation by subalgebras and, in particular, in RBF approximation, ridge function approximation, and neural network models.
+Recent papers (see [1, 4, 5, 6, 18]) improve the two-dimensional theory in connection with these directions.
+Moreover, based on the “levelling” algorithm, the ADI (alternating-direction-implicit) iteration method for numerical solution of differential equations is developed (see [38, 39, 40]).
+In the same paper, Diliberto and Straus stated that a similar algorithm will work for $n\geq 3$.
 However, G.
-Aumann in his work [3] has shown that while the proof in [6] was complete for case $n=2$, the generalization of the algorithm suggested for $n\geq 3$ was incorrect.
-A direct counterexample refuting the ”leveling algorithm” in $n\geq 3$ was given by V.A.
+Aumann in his work [7] has shown that while the proof in [11] was complete for case $n=2$, the generalization of the algorithm suggested for $n\geq 3$ was incorrect and proved a necessary and sufficient condition on the class of approximating functions that the “leveling algorithm” works.
+A direct counterexample refuting the “leveling algorithm” in $n\geq 3$ was given by V.A.
 Medvedev [29].
-In the paper [3] G.
-Aumann showed a necessary and sufficient condition on the class of approximating functions that the ”leveling algorithm” works.
-Another generalization approach for $n\geq 3$ to the problem starts with the paper [10] by M.
-Golomb, where he gave an important duality formula for the problem (1.1).
-Moreover, the works of G.
-Aumann and [17] by D.E.
+This example shows that the levelling algorithm, nor its small variations, can converge to a best approximation from the sum of three subspaces (see [34]).
+Another solution approach for $n\geq 3$ to the problem starts with the paper [16] by M.
+Golomb, where an important duality formula involving functionals defined on projection-cycles for the problem (1.1) was stated.
+However, later in [28] by D.E.
 Marshall and A.G.
-O’Farrell have shown that the proof of the duality given in [10] also had a gap.
-Future, based on the results of S.Ya.
+O’Farrell, it was shown that the proof of the formula given in [16] also has a gap.
+In [28], using the ergodic method, the lightning-bolt principle is also proven, which corresponds to the two-dimensional version of Golomb’s formula.
+It is worth remarking that the results obtained in the same paper in the two-dimensional case do not extend to sums such as $f(x)+g(y)+h(z)$ on subsets of $\mathbb{R}^{3}$.
+There is no way to generate the annihilators of a sum of three algebras using any analogue of the lightning bolts.
+This is related to the fact that ergodic theory is essentially limited to actions of groups that have subexponential growth.
+This shows that there are essentially new phenomena that arise with the sum of three subalgebras of $C(X)$, as opposed to two (see [32]).
+Future, with the help of the results of S.Ya.
 Khavinson [24] and K.G.
@@
 Golomb is valid in a stronger form.
-Note that for the two-dimensional case, under the condition $f_{xy}\geq 0$ T.J. Rivlin and R.J. Sibher [21] established that
+Although this formula itself involves finding a supremum over an infinite number of functionals and there is no closed formula to find the error explicitly, even with no additional conditions.
+The first finite formula for the two-dimensional case is established by T.J.Rivlin and R.J.Sibner in the paper [36], under the condition $f_{xy}\geq 0$,
 $E(f,[0,1]^{2})=\frac{1}{4}[f(1,1)-f(1,0)-f(0,1)+f(0,0)].$
 L.
-Flatto [7] gave a different proof of this formula in his work, where a generalization to the approximation of a $n$ variable function by $n-1$ variable functions was suggested under similar conditions.
-In case $n\geq 3$ (even with additional conditions), there is no finite algorithm for calculate the error $E(f,X)$.
-In this paper, under the conditions $\Delta_{x}\Delta_{y}f\geq 0$, $\Delta_{x}\Delta_{z}f\geq 0$, $\Delta_{y}\Delta_{z}f\geq 0$, we will prove a formula to calculate the error (1.1) for the three-variable continuous function on $[0,1]^{3}$.
+Flatto gave a different proof of the last formula in his work [12], where a general result for the approximation of an $n$ variable function by n1n-1 variable functions was suggested under similar conditions.
+Some other notable improvements in this direction have been made in [19] [20], where generalisations of the above formula in stair-like polygons have been proven.
+However, no finite formula has been proven for other domains, even for the simplest ones as the circle and triangle.
+In case $n\geq 3$ (even with additional conditions), there is no finite formula and converging algorithm to calculate the error $E(f,X)$.
+In the present paper, under conditions $\Delta_{x}\Delta_{y}f\geq 0$, $\Delta_{x}\Delta_{z}f\geq 0$, $\Delta_{y}\Delta_{z}f\geq 0$, we will prove a finite formula to calculate error (1.1) in $0,1]^{3}$ (See Theorem 2.2).
+The obtained results and methods proposed in this paper can be used in or extended to approximation by subalgebras and, in particular, in RBF approximation, ridge function approximation, and neural network models in the three-dimensional case.
+Moreover, the same ideas could lead to the construction of a three-dimensional levelling-type algorithm to find the best approximation functions, too.
+Such an algorithm would allow the construction of more efficient versions of existing numerical solution methods.
+As an example, the three-dimensional ADI iteration method itself proposed in [39] uses the two-dimensional levelling algorithm due to the lack of a converging three-dimensional alternative.
 1.2 Basic definitions and formulas
-Let’s give some definitions that will be used throughout this paper. Most definitions in this subsection are taken from the paper [12] (also can be found in the monograph [11]).
+We start with some definitions that will be used throughout this paper. Most of the definitions in this subsection are taken from the paper [23] (also can be found in the monograph [20]).
 Definition 1.1.
 A set of points $p=(x_{1},x_{2},...,x_{m}),\ x_{i}\in X$ where the set $X$ is the Cartesian product $X=X_{1}\times X_{2}\times\dots X_{n}$, is called a projection cycle if there exist some nonzero real numbers $\lambda_{1},\lambda_{2},..\lambda_{m}$ such that
-$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\forall j,0\leq j\leq n\text{ and }\forall\xi\in X_{j}$ (1.2)
+$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\ \forall\xi\in X_{j}\ \text{ and }\ \forall j,\ 0\leq j\leq n,$ (1.2)
 where $\pi_{j}(x)$ is the projection of $x$ into $X_{j}$ and $\xi$ is an arbitrary element of the set $X_{j}$.
-The definition of a projection cycle goes back to the M.Golomb [10]. The above-given definition is an equivalent form of Golomb’s definition taken from [12].
-Remark 1.1.
-For $n=3$, the left-hand side of the equation (1.2) becomes the sum over planes parallel to one of the coordinate planes.
-Thus, if a projection-cycle has a point on a plane parallel to one of the coordinate planes, then it has at least one other point with an opposite sign on the same plane.
+The definition of a projection cycle goes back to M.Golomb [16]. The above-given definition is an equivalent form of the Golomb’s definition taken from [23].
 Definition 1.2.
@@
 And $\lambda_{i}=\lambda(x_{i})$ is called the weight at point $x_{i}$.
-When we are given a weak projection cycle $p$ and a nonzero vector $\lambda$ satisfying the equation (1.3), by $(p,\lambda)$ we mean the projection cycle-vector based on the points of $p$ corresponding to nonzero components of $\lambda$.
+Remark 1.1.
+For $n=3$, the left-hand side of equation (1.2) becomes the sum over planes parallel to one of the coordinate planes.
+Thus, if a projection-cycle has a point on a plane parallel to one of the coordinate planes, then it has at least one other point with an opposite signed weight on the same plane.
 Definition 1.3.
 A projection cycle is called a minimal projection cycle if it has no proper subset that is a projection cycle.
-A projection cycle-vector $(p,\lambda)$ is called minimal if the set of points of $p$ corresponding to nonzero components of $\lambda$ is a minimal projection cycle.
+A projection cycle-vector $(p,\lambda)$ is called a minimal projection cycle-vector if the set of points of $p$ corresponding to nonzero components of $\lambda$ is a minimal projection cycle.
 Definition 1.4.
 A projection cycle-vector $(p,\lambda)$ is called orthonormal if
-$\sum|\lambda_{i}|=1$
-Definition 1.5.
-A set of points $p=(x_{1},x_{2},...,x_{m}),x_{i}\in X$ where the set $X$ is the Cartesian product $X=X_{1}\times X_{2}\times\dots X_{n}$, is called a weak projection cycle if there exist some real numbers $\lambda_{1},\lambda_{2},..\lambda_{m}$ such that, at least one $\lambda_{i}$ is nonzero and,
-$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\forall j,0\leq j\leq n\text{ and }\forall\xi\in X_{j}$ (1.3)
-where $\pi_{j}(x)$ is the projection of $x$ into $X_{j}$ and $\xi$ is an arbitrary element of the set $X_{j}$.
-The difference between a projection cycle and a weak projection cycle is that the second may contain points with zero weight.
-Thus, any projection-cycle is a weak projection cycle, but the converse is not true in general.
-For example, the set of points $\{(0,0),(0,1),(1,0),(1,1),(2,2)\}$ in $\mathbb{R}^{2}$ is a weak projection cycle but not a projection cycle.
-The concept of a minimal projection cycle and a weak projection cycle was defined by name loop and weak loop in [19] by K.G. Navada.
-The following theorem expresses Golomb’s formula (proposed in [10]), in revisited and improved form, from [12] (See: Theorem 2.8 in [12] or Theorem 3.37 in [11]).
+$\sum|\lambda_{i}|=1.$
+The concept of the minimal projection cycle was defined by name loop in [31] by K.G. Navada.
+The following theorem expresses Golomb’s formula (proposed in [16]), in a revisited and improved form, from [23] (See: Theorem 2.8 in [23] or Theorem 3.37 in [20]).
 Theorem 1.1 (Golomb’s Formula).
@@
 Remark 1.2.
-However, it is not a calculable formula; the formula (1.4) can be improved by comparing orthonormal sums over distinct projection-cycles under conditions like [21] or [7] used in their works to get a formula.
-In this paper, we will find a formula for the case $n=3$ and $X=[0,1]^{3}$.
-Remark 1.3.
-For simplicity, we use the term ”projection cycle-vector” instead of ”projection cycle, vector pair” used in [12] and [11].
+For simplicity, we use the term “projection cycle-vector” instead of “projection cycle, vector pair” used in [23] and [20].
+In what follows, the following notations will be used. As usual $\Delta_{(x,y,z)}f$ is defined as
+$\Delta_{(x,y,z)}f(u,v,w)=f(u+x,v+y,w+z)-f(u,v,w).$
+For simplicity, we denote
+$\Delta_{x}=\Delta_{(x,0,0)},\Delta_{y}=\Delta_{(0,y,0)},\Delta_{z}=\Delta_{(0,0,z)}.$
+For any finite sets $S$, the set $[S]$ denotes the set of all projection cycle-vectors, which, each point belongs to the set $S$ and $S]_{min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in[S]$.
 2 Three variable case
-Let $\Omega=[0,1]^{3}$ denote the unit cube and $f:\Omega\rightarrow\mathbb{R}$ be a given continuous function on $3$ variables. Consider the problem:
-$E(f,\Omega)=\inf||f(x,y,z)-\phi(x)-\psi(y)-\omega(z)||_{\infty}=?$
+Let $\Omega=[0,1]^{3}$ denote the unit cube, and let $f:\Omega\rightarrow\mathbb{R}$ be a given continuous function of three variables. Consider the problem:
+[…]
 where the infimum runs over all triplets of continuous functions $\phi(x),\psi(y),\omega(z)$ defined on the unit interval $[0,1]$.
-Let, $\Delta_{(x,y,z)}f$ be defined as
-$\Delta_{(x,y,z)}f(u,v,w)=f(u+x,v+y,w+z)-f(u,v,w)$
-For simplicity, let’s denote
-$\Delta_{x}=\Delta_{(x,0,0)},\Delta_{y}=\Delta_{(0,y,0)},\Delta_{z}=\Delta_{(0,0,z)}$
 Assume $f$ satisfies the following conditions at any point of $\Omega$
-$\begin{cases}\Delta_{x}\Delta_{y}f\geq 0\\ \Delta_{x}\Delta_{z}f\geq 0\\ \Delta_{y}\Delta_{z}f\geq 0\end{cases}$ (2.1)
+$\begin{cases}\Delta_{x}\Delta_{y}f\geq 0\\ \Delta_{x}\Delta_{z}f\geq 0\\ \Delta_{y}\Delta_{z}f\geq 0.\end{cases}$ (2.1)
 If $f$ has mixed derivatives with respect to any pair of variables at any point of $\Omega$, then the inequalities (2.1) are equivalent to
-$\begin{cases}f_{xy}\geq 0\\ f_{yz}\geq 0\\ f_{xz}\geq 0\end{cases}$ (2.2)
-Let, for any finite sets $S$, the set $X_{S}$ denote the set of all projection cycle-vectors, which, each point belongs to the set $S$ and $X_{S,min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in X_{S}$.
-We obtain the following result.
-Theorem 2.1.
-Let, $f\in C(\Omega)$ be a function satisfying the relations (2.1), then there exist a finite set $U=U_{f}$ such that,
-$E(f,\Omega)=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z},min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$ (2.3)
-where, $U_{x},U_{y},U_{z}$ are the projections of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively.
-Remark 2.1.
-For $n=2$, above-mentioned papers [6],[7], [11] gives similar teorems.
-But, while the set $U$ was the corners of the unit square, which is independent of $f$ in the case $n=2$, $U$ will be dependent on the function $f$ for $n=3$.
-Below, we will prove the Theorem $2.1$ and moreover will explicitly construct the set $U$.
-We start with applying Golomb’s Formula (1.4) and get:
+$\begin{cases}f_{xy}\geq 0\\ f_{yz}\geq 0\\ f_{xz}\geq 0.\end{cases}$ (2.2)
+We start by applying Golomb’s Formula (1.3) and get:
 $E(f,\Omega)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.3)
-where, the supremum runs over all minimal projection cycle-vectors $(p,\lambda)$ on the set $\Omega$.
-To proceed further, we will search the supremum (2.4) as
-$E(f,\Omega)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)$ (2.5)
-where the last supremum runs over all triples of finite sets $A,B,C\subset[0,1]$ Obviously, for any fixed $U$ and any projection cycle-vector $(p,\lambda)$, there exist at least one triple of sets $A,B,C\subset[0,1]$ with $U\subset A\times B\times C$ such that $(p,\lambda)\in X_{A\times B\times C}$.
-On the other hand choosing the sets $A,B,C$ finite ensures that the above maximum over $X_{A\times B\times C}$ is attained.
+where the supremum runs over all minimal projection cycle-vectors $(p,\lambda)$ on the set $\Omega$. To proceed further, we will search the supremum (2.3) as
+$E(f,\Omega)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\sup_{A,B,C}\left({\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right),$ (2.4)
+where the last supremum runs over all triples of finite sets $A,B,C\subset[0,1]$.
+Obviously, for any projection cycle-vector $(p,\lambda)$, there exists at least one triple of sets $A,B,C\subset[0,1]$ such that $(p,\lambda)\in[A\times B\times C]$.
+On the other hand, choosing the sets $A,B,C$ finite ensures that the above maximum over $A\times B\times C]$ is attained.
 Hence, the relation (2.4) is well-defined.
-Before proving the Theorem $2.1$ we have to prove some facts.
-Later, we will see that to prove the Theorem 2.1 for the class of functions satisfying the conditions (2.2) is equivalent to proving it for functions satisfying the following stricter inequalities
-$\begin{cases}\Delta_{x}\Delta_{y}f>0\\ \Delta_{x}\Delta_{z}f>0\\ \Delta_{y}\Delta_{z}f>0\end{cases}$ (2.6)
-Thus, below here we will work with functions satisfying (2.6).
+Before proving our main results, we have to prove some facts. From now on will proceed with the following stricter version of the inequalities (2.1)
+$\begin{cases}\Delta_{x}\Delta_{y}f>0\\ \Delta_{x}\Delta_{z}f>0\\ \Delta_{y}\Delta_{z}f>0.\end{cases}$ (2.5)
+Later, we will see that we can pass back to the original condition (2.1).
 2.1 Optimal Projection Cycle-Vectors on a Finite Set
-First, we will investigate some properties of projection cycle-vectors that give the following maximum
-$\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.7)
-where $A,B,C\subset[0,1]$ are arbitrary chosen sets with $\{0,1\}\subset A\cap B\cap C$.
-Let’s give some further definitions and auxiliary facts.
+We start by investigating properties of projection cycle-vectors that give the following maximum
+$\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|},$ (2.6)
+where $A,B,C\subset[0,1]$ are the sets chosen arbitrary with $\{0,1\}\subset A\cap B\cap C$ and $f\in C(\Omega)$ is a function satisfying the conditions (2.1) or (2.5).
+Let us give some further definitions and auxiliary facts.
 Definition 2.1.
 Let’s call two points $(\xi_{1},\gamma_{1},\zeta_{1})$ and $(\xi_{2},\gamma_{2},\zeta_{2})$ of the space $\mathbb{R}^{3}$ well-ordered if either
-$\xi_{1}\geq\xi_{2},\gamma_{1}\geq\gamma_{2},\zeta_{1}\geq\zeta_{2}\text{ or }\xi_{1}\leq\xi_{2},\gamma_{1}\leq\gamma_{2},\zeta_{1}\leq\zeta_{2}.$
+[…]
 In other words, two points are called well-ordered if all components of one of the points are not smaller than the corresponding coordinates of the other point.
-If two points$(\xi_{1},\gamma_{1},\zeta_{1})$ and $(\xi_{2},\gamma_{2},\zeta_{2})$ are not well-ordered, then, without loss of generality, we can assume that:
+[…]
 $\xi_{1}\geq\xi_{2},\ \gamma_{1}>\gamma_{2},\ \zeta_{1}<\zeta_{2}.$
 All other versions can be transferred into this, either by changing the names of variables or the names of points.
-Now we present an important lemma, which is a direct consequence of (2.6).
+Now we present an important lemma, which is a direct consequence of the inequalities (2.5).
 Lemma 2.1.
@@
 $f(\xi_{1},\gamma_{1},\zeta_{1})+f(\xi_{2},\gamma_{2},\zeta_{2})\leq f\left((\xi_{1},\xi_{2}),(\gamma_{1},\gamma_{2})(\zeta_{1},\zeta_{2})\right)+f\left([\xi_{1},\xi_{2}],[\gamma_{1},\gamma_{2}],[\zeta_{1},\zeta_{2})]\right)$ (2.7)
-holds, where we denote $(a,b)=\min(a,b)$ and $[a,b]=\max(a,b)$. The equality happens iff $x_{1}$ and $x_{2}$ are well-ordered.
+holds, where we denote $(a,b)=\min(a,b)$ and $a,b]=\max(a,b)$. The equality happens iff $x_{1}$ and $x_{2}$ are well-ordered.
 Proof.
-If $x_{1}$ and $x_{2}$ are well-ordered, then the expressions on the left and right-hand sides become the same, and the equality holds.
+If $x_{1}$ and $x_{2}$ are well-ordered, then the expressions on the left- and right-hand sides become the same, and the equality holds.
 If $x_{1}$ and $x_{2}$ are not well-ordered, without loss of generality we can assume that $\xi_{1}\geq\xi_{2}$, $\gamma_{1}>\gamma_{2}$ and $\zeta_{1}<\zeta_{2}$. Thus, we have
@@
 $f(x_{1}^{\prime})+f(x_{2}^{\prime})-f(x_{1})-f(x_{2})$
-$=\left(f(x_{2}^{\prime})+f(x_{2}^{\prime\prime})-f(x_{1}^{\prime\prime})-f(x_{2})\right)+\left(\left(f(x_{1}^{\prime\prime})+f(x_{1}^{\prime})-f(x_{2}^{\prime\prime})-f(x_{1})\right)\right)>0$
+$\left(f(x_{2}^{\prime})+f(x_{2}^{\prime\prime})-f(x_{1}^{\prime\prime})-f(x_{2})\right)+\left(\left(f(x_{1}^{\prime\prime})+f(x_{1}^{\prime})-f(x_{2}^{\prime\prime})-f(x_{1})\right)\right)>0,$
 which, directly follows from (2.5)(See: Figure 1). ∎
-$x$$z$$y$$x_{2}^{\prime\prime}$$x_{1}^{\prime\prime}$$x_{2}^{\prime}$$x_{2}$$x_{1}$$x_{1}^{\prime}$
+[…]
 Figure 1: Geometric interpretation of the proof of Lemma 2.1
 Definition 2.2.
-A projection cycle-vector $(p^{0},\lambda^{0})\in X_{S}$ is called optimal for the function $f$ on the finite set $S$ if
-$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
-The set of all optimal projection cycle-vectors for $f$ on $S$ is denoted by $X_{S,f}$.
-Obviously, for any $(p,\lambda)\in X_{S,f}$ we have
-$\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq 0$ (2.9)
-as otherwise, taking $(p,-\lambda)$ will contradict the optimality of $(p,\lambda)$.
+A projection cycle-vector $(p^{0},\lambda^{0})\in[S]$ is called optimal for the function $f$ on the finite set $S$ if
+$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
+The set of all optimal projection cycle-vectors for $f$ on $S$ is denoted by $S]_{f}$.
+Obviously, for any $(p,\lambda)\in[S]_{f}$ we have
+$\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq 0,$ (2.8)
+as otherwise, taking $(p,-\lambda)$ will contradict the optimality of $(p,\lambda)$. And, moreover since $[S]$ is finite, $S]_{f}$ is always non-empty.
 Definition 2.3.
 The sum of two projection cycles-vectors is a projection cycle-vector on the union of the points of the given projection cycle-vectors, with weight at each point equal to the sum of the weights of the given projection cycle-vectors on that point.
-We can express the above definition by formulas as follows: $(p^{1},\lambda^{1})+(p^{2},\lambda^{2})=(p,\lambda)$ where, $p=p^{1}\cup p^{2}$ and
-$\lambda(x)=\begin{cases}\lambda^{1}(x),\hskip 45.0ptx\in p^{1}\backslash p^{2}\\ \lambda^{1}(x)+\lambda^{2}(x),\hskip 8.0ptx\in p^{1}\cap p^{2}\\ \lambda^{2}(x),\hskip 43.0ptx\in p^{2}\backslash p^{1}\\ \end{cases}$ (2.10)
-for any two (weak) projection cycle vectors $(p^{1},\lambda^{1})$ and $(p^{2},\lambda^{2})$.
-Clearly,if both $(p^{1},\lambda^{1})\in X_{S}$ and $(p^{2},\lambda^{2})\in X_{S}$ for some finite set $S$ and function $f$, then
-$(p^{1},\lambda^{1})+(p^{2},\lambda^{2})\in X_{S}$
-since $p^{1}\subset S$ and $p^{2}\subset S$ is equivalent to $p^{1}\cup p^{2}\subset S$. Moreover, we have
-$\sum_{x_{i}\in p}\lambda_{i}f(x_{i})=\sum_{x_{j}\in p^{1}}\lambda^{1}_{j}f(x_{j})+\sum_{x_{k}\in p^{2}}\lambda^{2}_{k}f(x_{k})$
+If we encounter any points with a weight of zero, we would ignore them.
+We can express the above definition by formulas as follows:
+$(p^{1},\lambda^{1})+(p^{2},\lambda^{2})=(p,\lambda),$
+where $\pp=\left(p^{1}\cup p^{2}\right)\backslash p^{*}$, $p^{*}=\{q:q\in p^{1}\cap p^{2},\lambda^{1}(q)+\lambda^{2}(q)=0\}$ and
+[…]
+for any two projection cycle-vectors $(p^{1},\lambda^{1})$ and $(p^{2},\lambda^{2})$.
+Clearly, if both $(p^{1},\lambda^{1})\in[S]$ and $(p^{2},\lambda^{2})\in[S]$ for some finite set $S$, then
+$(p^{1},\lambda^{1})+(p^{2},\lambda^{2})\in[S]$
+since $p^{1}\subset S$ and $p^{2}\subset S$ is equivalent to $p^{1}\cup p^{2}\subset S$. Moreover for any function $f$ defined on $0,1]^{3}$, we have
+$\sum_{x_{i}\in p}\lambda_{i}f(x_{i})=\sum_{x_{j}\in p^{1}}\lambda^{1}_{j}f(x_{j})+\sum_{x_{k}\in p^{2}}\lambda^{2}_{k}f(x_{k}).$
+In fact, we will never encounter a situation of adding two projection cycle-vectors of the form $(p,\lambda)$ and $(p,-\lambda)$.
+However, for the sake of correctness, we can still denote this sum as a vanishing projection cycle-vector.
+That is a projection cycle, with no points and no weights, and will not be considered an element of $[S]$, for any set $S$.
+The following lemma demonstrates our way of constructing a “more optimal” projection-cycle vector from the given “non-optimal” one.
 Lemma 2.2.
-Let $(p^{1},\lambda^{1})$, and $(p^{2},\lambda_{\varepsilon})$ be two projection cycle vectors, where, $\lambda_{\varepsilon}$ is a vector of length $m$ such that each component is $\varepsilon$ or $-\varepsilon$ for some positive real number $\varepsilon$.
-Let $k$ be the number of common points of $p^{1}$ and $p^{2}$, such that the corresponding components of the vectors $\lambda$ and $\lambda_{\varepsilon}$ have opposite signs.
+Let $(p^{1},\lambda^{1})$ and $(p^{2},\lambda_{\varepsilon})$ be two projection cycle vectors, where $\lambda_{\varepsilon}$ is a vector of length $m$ such that each component is $\varepsilon$ or $\varepsilon$ for some positive real number $\varepsilon$.
+Let $k$ be the number of common points of $p^{1}$ and $p^{2}$, such that the corresponding components of the vectors $\lambda^{1}$ and $\lambda_{\varepsilon}$ have opposite signs.
 If $k\geq\frac{m}{2}$, then for
@@
 we have
-$\sum_{x\in p}|\lambda(x)|\leq\sum_{x\in p^{1}}|\lambda^{1}(x)|$ (2.11)
-where, $(p,\lambda)=(p^{1},\lambda^{1})+(p^{2},\lambda_{\varepsilon})$.
+$\sum_{x\in p}|\lambda(x)|\leq\sum_{x\in p^{1}}|\lambda^{1}(x)|,$ (2.10)
+where $(p,\lambda)=(p^{1},\lambda^{1})+(p^{2},\lambda_{\varepsilon})$.
 Proof.
-By the definition (2.10) we have,
-$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|=\sum_{x\in p^{1}\cap p^{2}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon$
-The first summand can be written as,
+By the equation (2.9) we have
+$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|=\sum_{x\in p^{1}\cap p^{2}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon.$
+The first summand can be written as
 $\sum_{x\in p^{1}\cap p^{2}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)=$
-$=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)$
+$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right).$
 By the triangle inequality, we have
-$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)\leq\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}|\lambda_{\varepsilon}(x)|=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon$
+$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)\leq\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}|\lambda_{\varepsilon}(x)|=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon.$
 On the other hand, for
@@
 we have
-$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)=-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon$
+$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)=-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon.$
 Thus, we have
-$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|\leq-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon\leq 0$
-Where the last inequality follows from the fact that the number of summands with a negative sign is $k$, the number of all summands is $m$ and $k\geq\frac{m}{2}$
+$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|\leq-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon\leq 0.$
+Where the last inequality follows from the fact that the number of summands with a negative sign is $k$, the number of all summands is $m$, and $k\geq\frac{m}{2}$
 The next theorem deals with the geometric structure of optimal projection cycles.
@@
   2. (b)
-    Let’s call a plane on $\mathbb{R}^{3}$ an interior plane of the unit cube if it is parallel to one of the faces of the unit cube and passes through an interior point of the cube.
+    Let us call a plane on $\mathbb{R}^{3}$ an interior plane of the unit cube if it is parallel to one of the faces of the unit cube and passes through an interior point of the cube.
 If $p$ contains a well-ordered pair of points $x_{i},x_{j}$ lying on the same interior plane, then, $\lambda_{i},\lambda_{j}>0$,
   3. (c)
-    If $\lambda_{i}<0$, then the point $x_{i}$ belongs to one of the edges of the unit cube not containing any of $(0,0,0)$ and $(1,1,1)$,
+    If $\lambda_{i}<0$, then the point $x_{i}$ belongs to one of the edges of the unit cube that is not adjacent to any of $(0,0,0)$ and $(1,1,1)$,
   4. (d)
@@
   5. (e)
-    $(0,0,0),(1,1,1)\in p$ and, both has positive weights.
+    Both of the points $(0,0,0)$ and $(1,1,1)$ belong to $p$ and, both have positive weights.
 Proof.
-(a) Assume the contrary: Let, $(p,\lambda)=((x_{0},x_{1},...,x_{m}),(\lambda_{0},\lambda_{1},...,\lambda_{m}))\in X_{S,f}$ contain a non well-ordered pair of points $x_{k},x_{j}$ with $\lambda_{k}>0,\lambda_{j}>0$.
-Let the coordinates of the points $x_{k},x_{j}$ be $(\xi_{k},\gamma_{k},\zeta_{k})$ and $(\xi_{j},\gamma_{j},\zeta_{j})$ respectively.
+(a) Assume the contrary: Let $(p,\lambda)=((x_{0},x_{1},...,x_{m}),(\lambda_{0},\lambda_{1},...,\lambda_{m}))\in[S]_{f}$ contain a non well-ordered pair of points $x_{k},x_{j}$ with $\lambda_{k}>0,\lambda_{j}>0$.
+Let the coordinates of the points $x_{k},x_{j}$ be $(\xi_{k},\gamma_{k},\zeta_{k})$ and $(\xi_{j},\gamma_{j},\zeta_{j})$, respectively.
 Since $x_{k},x_{j}$ are not well-ordered, without loss of generality we can assume that $\xi_{k}\geq\xi_{j}$, $\gamma_{k}>\gamma_{j}$ and $\zeta_{k}<\zeta_{j}$.
-Denote the points $(\xi_{k},\gamma_{k},\zeta_{j})$ and $(\xi_{j},\gamma_{j},\zeta_{k})$ by $x_{k}^{\prime}$ and $x_{j}^{\prime}$ respectively.
-Clearly, $x_{k}^{\prime},x_{j}^{\prime}\in S$ and the projection cycle-vector $(p_{0},\lambda_{\varepsilon})\in X_{S}$ with $p_{0}=(x_{k},x_{j},x_{k}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,-\varepsilon,\varepsilon,\varepsilon)$ is well defined for any $\varepsilon>0$.
+Denote the points $(\xi_{k},\gamma_{k},\zeta_{j})$ and $(\xi_{j},\gamma_{j},\zeta_{k})$ by $x_{k}^{\prime}$ and $x_{j}^{\prime}$, respectively.
+Clearly, $x_{k}^{\prime},x_{j}^{\prime}\in S$ and the projection cycle-vector $(p_{0},\lambda_{\varepsilon})\in[S]$ with $p_{0}=(x_{k},x_{j},x_{k}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,-\varepsilon,\varepsilon,\varepsilon)$ is well defined for any $\varepsilon>0$.
 Let, $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$.
-Since $\lambda(x_{k})$ and $\lambda(x_{j})$ are positive, when $\varepsilon>0$ is small enough by Lemma $2.1$ we have $0\leq\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$.
-On the other hand,
-$\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})=\varepsilon(f(x^{\prime}_{k})+f(x^{\prime}_{j})-f(x_{k})-f(x_{j}))>0$
+Since $\lambda(x_{k})$ and $\lambda(x_{j})$ are positive, when $\varepsilon>0$ is small enough by Lemma 2.2 we have $0\leq\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$. On the other hand,
+$\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})=\varepsilon(f(x^{\prime}_{k})+f(x^{\prime}_{j})-f(x_{k})-f(x_{j}))>0,$
 where the last inequality follows by Lemma 2.1 applied to the points $x_{k}$ and $x_{j}$ and the positivity of $\varepsilon$.
-Now since, $\sum\lambda^{\prime}_{i}f(x_{i})>\sum\lambda_{i}f(x_{i})\geq 0$ and $0<\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$, we have
-$\frac{\sum\lambda^{\prime}_{i}f(x_{i})}{\sum|\lambda^{\prime}_{i}|}-\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\frac{\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}>0$
-Which contradicts to the optimality of $p,\lambda)$.
+Now, since $\sum\lambda^{\prime}_{i}f(x_{i})>\sum\lambda_{i}f(x_{i})\geq 0$ and $0<\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$, we have
+$\frac{\sum\lambda^{\prime}_{i}f(x_{i})}{\sum|\lambda^{\prime}_{i}|}-\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\frac{\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}>0.$
+This contradicts the optimality of $(p,\lambda)$.
 (b) Let $(p,\lambda)=((x_{0},x_{1},...,x_{m}),(\lambda_{0},\lambda_{1},...,\lambda_{m}))$ be an optimal projection cycle-vector for $f$ on the set $S$ that contains a well-ordered pair of points $x_{i},x_{j}$ from the same interior plane.
-Without loss of generality we can assume that, $x_{i}$ and $x_{j}$ are lying on the same plane paralel to the $xOy$ plane and thus, we can write, $x_{i}=(\xi_{i},\gamma_{i},\zeta)$ and $x_{j}=(\xi_{j},\gamma_{j},\zeta)$ for some numbers $\xi_{i},\xi_{j},\gamma_{i},\gamma_{j}\in[0,1]$ and some $\zeta\in(0,1)$.
+Without loss of generality, we can assume that $x_{i}$ and $x_{j}$ are lying on the same plane parallel to the $xOy$ plane, and thus we can write $x_{i}=(\xi_{i},\gamma_{i},\zeta)$ and $x_{j}=(\xi_{j},\gamma_{j},\zeta)$ for some numbers $\xi_{i},\xi_{j},\gamma_{i},\gamma_{j}\in[0,1]$ and some $\zeta\in(0,1)$.
 Assume $\xi_{i}\leq\xi_{j}$, then, since $x_{i},x_{j}$ are well-ordered we have $\gamma_{i}\leq\gamma_{j}$. If both $\lambda_{i},\lambda_{j}$ are not positive, have three cases:
-  1. Case 1:
+  1. 1.
     $\lambda_{i}>0>\lambda_{j}$
-    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},0)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},0)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$
-  2. Case 2:
+    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},0)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},0)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$;
+  2. 2.
     $\lambda_{i}<0<\lambda_{j}$
-    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},1)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},1)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,-\varepsilon,-\varepsilon,\varepsilon)$
-  3. Case 3:
+    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},1)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},1)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,-\varepsilon,-\varepsilon,\varepsilon)$;
+  3. 3.
     $\lambda_{i},\lambda_{j}<0$
-    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{j},\zeta)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{i},\zeta)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$
-$x$$z$$y$$x_{i}^{\prime}$$x_{i}$$x_{j}^{\prime}$$x_{j}$
+    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{j},\zeta)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{i},\zeta)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$.
+[…]
 Figure 2: Case 1: We simply replace,the pair $x_{i},x_{j}$ with $x^{\prime}_{i},x^{\prime}_{j}$
 In each case we have chosen $\lambda_{\varepsilon}$ in a way that, at both of the points $x_{i}$ and $x_{j}$ the projection cycle-vectors $(p,\lambda)$ and $(p_{0},\lambda_{\varepsilon})$ have weights with opposite signs.
-Thus, by the Lemma $2.2$, for $\varepsilon>0$ small enough, we have
+Thus, by the Lemma 2.2, for $\varepsilon>0$ small enough, we have
 $0<\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|,$
@@
 Consider the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$.
-By the definition of the projection cycle each of these planes contains at least one point of $p$ with positive weight.
+By the definition of the projection cycle, each of these planes contains at least one point of $p$ with positive weight.
 Let these points be $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j}),x_{l}=(\xi_{l},\gamma_{i},\zeta_{l}),x_{k}=(\xi_{k},\gamma_{k},\zeta_{i})$.
 If any of these positive-weighted points makes a well-ordered pair with $x_{i}$, then the projection cycle-vector $(p,\lambda)$ cannot be optimal by part $(b)$.
-$x$$z$$y$$x_{i}$
+[…]
 Figure 3: The possible locations of the points that balances $x_{i}$
 due to part $(b)$.
-Assume both $10<\xi<$ 1 and $10<\gamma<$ 1.
-By the part $(b)$ we have $(x_{i},x_{j})$ is not a well-ordered pair,thus, either $\gamma_{i}<\gamma_{j},\zeta_{i}>\zeta_{j}$ or $\gamma_{i}>\gamma_{j},\zeta_{i}<\zeta_{j}$.
+Assume both $0<\xi<1$ and $0<\gamma<1$.
+By the part $(b)$ we have that $(x_{i},x_{j})$ is not a well-ordered pair,thus, either $\gamma_{i}<\gamma_{j},\zeta_{i}>\zeta_{j}$ or $\gamma_{i}>\gamma_{j},\zeta_{i}<\zeta_{j}$.
 Similarly, $(x_{i},x_{l})$ is not a well-ordered pair, thus, either $\xi_{i}<\xi_{l},\zeta_{i}>\zeta_{l}$ or $\xi_{i}>\xi_{l},\zeta_{i}<\zeta_{l}$.
-It’s easy to verify that in each of the four possible cases, $x_{j},x_{l}$ cannot be a well-ordered pair.
-By part $(a)$ if $(p,\lambda)$ is optimal then any pair of points $x_{j},x_{l},x_{k}$ should be well-ordered.
+It is easy to verify that in each of the four possible cases, $x_{j},x_{l}$ cannot be a well-ordered pair.
+By part $(a)$ if $(p,\lambda)$ is optimal, then any pair of points $x_{j},x_{l},x_{k}$ should be well-ordered.
 Thus, $(p,\lambda)$ cannot be optimal, and we get a contradiction.
-Since, we choose $\xi,\gamma$ arbitrarily from $\xi,\gamma,\zeta$, this contradiction works if at least two of the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$ are inner planes, which happens when $x_{i}$ is not on the edges of the unit cube.
+Since we choose $\xi,\gamma$ arbitrarily from $\xi,\gamma,\zeta$, this contradiction works if at least two of the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$ are inner planes, which happens when $x_{i}$ is not on the edges of the unit cube.
 Hence, $x_{i}$ belongs to one of the edges of the unit cube.
 Now, if, $x_{i}$ and $(0,0,0)$ are on the same edge, we have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$.
-Assume, $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist an other point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}>0$.
-Moreover, since, $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered.
+Assume that $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist another point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}>0$.
+Moreover, since $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered.
 But this contradicts the optimality of $(p,\lambda)$, by part $(b)$.
@@
 Using similar arguments, it can be proven that $x_{i}$ and $(1,1,1)$ cannot be on the same edge.
-Thus, $x_{i}$ belongs to one of the edges of the cube, not containing any of the vertices $(0,0,0)$ and $(1,1,1)$. Let’s name these edges $l_{1},...,l_{6}$ as following
-$l_{1}=\{(\xi,0,1),0\leq\xi\leq 1\}$
-$l_{2}=\{(1,0,\zeta),0\leq\zeta\leq 1\}$
-$l_{3}=\{(1,\gamma,0),0\leq\gamma\leq 1\}$
-$l_{4}=\{(\xi,1,0),0\leq\xi\leq 1\}$
-$l_{5}=\{(0,1,\zeta),0\leq\zeta\leq 1\}$
-$l_{6}=\{(0,\gamma,1),0\leq\gamma\leq 1\}$
+Thus, $x_{i}$ belongs to one of the edges of the cube, which does not contain any of the vertices $(0,0,0)$ and $(1,1,1)$. Let’s name these edges $l_{1},...,l_{6}$ as following
+$l_{1}=\{(\xi,0,1),0\leq\xi\leq 1\},$
+$l_{2}=\{(1,0,\zeta),0\leq\zeta\leq 1\},$
+$l_{3}=\{(1,\gamma,0),0\leq\gamma\leq 1\},$
+$l_{4}=\{(\xi,1,0),0\leq\xi\leq 1\},$
+$l_{5}=\{(0,1,\zeta),0\leq\zeta\leq 1\},$
+$l_{6}=\{(0,\gamma,1),0\leq\gamma\leq 1\}.$
 (d) Assume the contrary, let $x_{i}\in p$ be a point on one of the edges of the unit cube with $\lambda_{i}>0$.
 We have three cases: Case 1: $x_{i}$ is on the same edge with one of the vertices $(0,0,0)$ or $(1,1,1)$.
-If, $x_{i}$ and $(0,0,0)$ are on the same edge, have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$.
-Assume, $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist an other point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}<0$.
-Moreover, since, $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered.
+If $x_{i}$ and $(0,0,0)$ are on the same edge, have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$.
+Assume $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exists another point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}<0$.
+Moreover, since $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered.
 But this contradicts the optimality of $(p,\lambda)$, by part $(b)$.
@@
 To make the sum of weights on the plane $z=1$ equal to zero, there should exist a point $x_{j}$ of $p$ with a negative weight on $z=1$.
-By the part $(c)$ the point $x_{j}$ belongs to at least one of the edges $l_{1}$ or $l_{6}$.
+By part $(c)$ the point $x_{j}$ belongs to at least one of the edges $l_{1}$ or $l_{6}$.
 Thus, we have two sub-cases:
-Case 2.1: $x_{j}=(\xi_{j},0,1)$ for some $0\leq\xi_{j}\leq 1$ Clearly, $\xi_{j}\neq\xi$. We should consider two sub-cases:
+Case 2.1: $x_{j}=(\xi_{j},0,1)$ for some $0\leq\xi_{j}\leq 1$. Clearly, $\xi_{j}\neq\xi$. We should consider two sub-cases:
 Case 2.1.1: $\xi_{j}<\xi$
-Consider the projection cycle $(p,\lambda_{\varepsilon})$, where $p_{0}=(x_{i},x_{j},(\xi,1,1),(\xi_{j},1,1))$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$
+Consider the projection cycle $(p,\lambda_{\varepsilon})$, where $p_{0}=(x_{i},x_{j},(\xi,1,1),(\xi_{j},1,1))$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$;
 Case 2.1.2: $\xi<\xi_{j}$
 Consider the projection cycle $(p_{0},\lambda_{\varepsilon})$, where $p_{0}=(x_{i},x_{j},(\xi,0,0),(\xi_{j},0,0))$ and $\lambda_{\varepsilon}=(\varepsilon,-\varepsilon,-\varepsilon,\varepsilon)$.
-By a similar way to the part $(a)$ and part $(b)$, taking $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$ we get a contradiction to the optimality of $(p,\lambda)$.
-Case 2.2: $x_{j}=(0,\gamma,1)$ for some $0<\gamma\leq 1$ Consider the plane, $y=\gamma$.
+Similarly to part $(a)$ and part $(b)$, taking $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$ we get a contradiction to the optimality of $(p,\lambda)$.
+Case 2.2: $x_{j}=(0,\gamma,1)$ for some $0<\gamma\leq 1$.
+Consider the plane, $y=\gamma$.
 To make the sum of weights on the plane $y=\gamma$ equal to zero, there should exist a point $x_{k}$ of $p$ with a positive weight on $y=\gamma$.
-Let $x_{k}=(\xi_{k},\gamma,\zeta)$.If $\zeta=1$, the points, $x_{j}$ and $x_{k}$ belonging to the inner plane $y=\gamma$ becomes well-ordered and since $x_{j}$ and $x_{k}$ have opposite weights we get a contradiction by part $(b)$.
+Let $x_{k}=(\xi_{k},\gamma,\zeta)$.
+If $\zeta=1$, the points $x_{j}$ and $x_{k}$ belonging to the inner plane $y=\gamma$ becomes well-ordered and since $x_{j}$ and $x_{k}$ have opposite weights we get a contradiction by part $(b)$.
 If, $\zeta<1$, the points, $x_{i}$ and $x_{k}$ becomes non well-ordered.
 Since both of $x_{i}$ and $x_{k}$ have positive weights, we get a contradiction by the part $(a)$.
-Since, we get a contradiction in each case, the proof of part $(d)$ is completed.
-(e) Consider the point $x_{i}=(\xi_{i},\gamma_{i},\zeta_{i})\in p$ with the minimal coordinates among positive weighted points of $p$.
-(Since, the positive weighted points are well-ordered pairwise, the point with the minimal sum of coordinates has minimal coordinates.)
+Since we get a contradiction in each case, the proof of part $(d)$ is completed.
+(e) Consider the point $x_{i}=(\xi_{i},\gamma_{i},\zeta_{i})\in p$ with minimal coordinates among the positive weighted points of $p$.
+(Since the positive weighted points are well-ordered pairwise, the point with the minimal sum of coordinates has minimal coordinates.)
 If $x_{i}=(0,0,0)$, we are done.
 Otherwise, since $x_{i}$ is chosen to be minimal, it cannot be any other vertex of the unit cube, and thus there is at least one inner plane passing through $x_{i}$.
-Without loss of generality, assume this plane has equation $x=\xi_{i}$.
+Without loss of generality, assume that this plane has equation $x=\xi_{i}$.
 Then, on the plane $x=\xi_{i}$ there is another point $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j})\in p$ such that $\lambda_{j}<0$.
-By part $(b)$ we know that $x_{i}$ and $x_{j}$ cannot be well-ordered and hence, either $\gamma_{j}<\gamma_{i}$ or $\zeta_{j}<\zeta_{i}$ but in the first case on the plane $y=\gamma_{j}$ and in the second case on the plane $z=\zeta_{j}$ there is a positive weighted point to make the sum over that plane equal to zero.
-But this contradicts to the minimality of $x_{i}$ and hence, $x_{i}=(0,0,0)$.
+By part $(b)$ we know that $x_{i}$ and $x_{j}$ cannot be well-ordered, and hence, either $\gamma_{j}<\gamma_{i}$ or $\zeta_{j}<\zeta_{i}$ but in the first case on the plane $y=\gamma_{j}$ and in the second case on the plane $z=\zeta_{j}$ there is a positive weighted point to make the sum over that plane equal to zero.
+But this contradicts to the minimality of $x_{i}$ and, hence, $x_{i}=(0,0,0)$.
 Similarly, we can prove that the positive weighted point with maximal coordinates is $(1,1,1)$.
-$x$$z$$y$
+xxzzyy
 Figure 4: Geometric interpretation: Theorem 2.2
-Let $X^{\prime}_{S}$ be the set of projection cycle-vectors $(p,\lambda)\in X_{S}$, such that the propositions $(a)-(e)$ of the Theorem 2.2 holds for $(p,\lambda)$ and let $X^{\prime}_{S,min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in X^{\prime}_{S}$.
-The following lemmas hold.
+From now on, $S]^{(a)-(e)}$ will denote the set of projection cycle-vectors $(p,\lambda)\in[S]$, such that the propositions (a)(e)(a)-(e) of the Theorem 2.2 holds for $(p,\lambda)$ and let $S]^{(a)-(e)}_{min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in[S]^{(a)-(e)}$.
+The following lemmas will later help us to prove our main results for functions satisfying the inequalities (2.1), instead of the stricter version (2.5).
 Lemma 2.3.
@@
 Proof.
-(See: Lemma 4 and Lemma 5 in [19] or Lemma 3.32 in [11]). ∎
+(See: [31, Lemma 4 and Lemma 5] or [20, Lemma 3.32]). ∎
 Lemma 2.4.
-For any projection cycle $(p,\lambda)\in X_{S}$ and for any function $f\in C(\Omega)$, there exist a projection cycle $(p^{0},\lambda^{0})\in X_{S,min}$ such that $p^{0}\subset p$ and the following relations hold
+For any projection cycle $(p,\lambda)\in[S]$ and for any function $f\in C(\Omega)$, there exist a projection cycle $(p^{0},\lambda^{0})\in[S]_{min}$ such that $p^{0}\subset p$ and the following relations hold
   1. 1.
-    $\lambda^{0}(x)\lambda(x)>0,\forall x\in p^{0}$
+    $\lambda^{0}(x)\lambda(x)>0,\ \forall x\in p^{0},$
   2. 2.
-    $\frac{\sum_{x_{j}\in p^{0}}\lambda^{0}_{j}f(x_{j})}{\sum|\lambda^{0}_{j}|}\geq\frac{\sum_{x_{i}\in p}\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
+    $\frac{\sum_{x_{j}\in p^{0}}\lambda^{0}_{j}f(x_{j})}{\sum|\lambda^{0}_{j}|}\geq\frac{\sum_{x_{i}\in p}\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
 Proof.
-See [19, Theorem 2]. ∎
-The Lemma 2.3 implies that, for any function $f$,
-$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
+See [31, Theorem 2]. ∎
+Lemma 2.4 implies that, for any function $f$,
+$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
 and
-$\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
+$\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
 Lemma 2.5.
-For any set $S=S_{x}\times S_{y}\times S_{z}$ where, $S_{x},S_{y},S_{z}\subset[0,1]$ are some finite sets with $\{0,1\}\subset S_{x}\cap S_{y}\cap S_{z}$ and for any function $f\in C(\Omega)$ satisfying the inequalities (2.1), the intersection $X^{\prime}_{S}\cap X_{S,f}$ is non-empty.
+For any set $S=S_{x}\times S_{y}\times S_{z}$, where $S_{x},S_{y},S_{z}\subset[0,1]$ are some finite sets with $\{0,1\}\subset S_{x}\cap S_{y}\cap S_{z}$ and for any function $f\in C(\Omega)$ satisfying the inequalities (2.1), the intersection $S]^{(a)-(e)}\cap[S]_{f}$ is non-empty.
 Proof.
 Let’s define
-$f_{n}(x,y,z)=f(x,y,z)+\frac{1}{n}(xy+xz+yz),\forall n\in\mathbb{N}$
+$f_{n}(x,y,z)=f(x,y,z)+\frac{1}{n}(xy+xz+yz),\ \forall n\in\mathbb{N}.$
 Clearly, the function $f_{n}$ satisfies the inequalities (2.5) and therefore it follows from Theorem 2.2 and Lemma 2.4 that,
-$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}$
-Here, we can change the order of the limit and the maximum as $X_{S,min}$ is a finite set by the Lemma 2.3 (considering vectors up to multiplication by a constant) and hence,
-$\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}$
+$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}.$
+Here, we can change the order of the limit and the maximum, as $S]_{min}$ is a finite set by Lemma 2.3 (considering vectors up to multiplication by a constant), and hence,
+$\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}.$
 By the Theorem 2.2 and Lemma 2.4,
-$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}$
+$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}.$
 Changing the order of the limit and maximum once more and using the Lemma 2.3 gives the desired result:
-$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}$
-Now let’s choose $(p^{0},\lambda^{0})\in X^{\prime}_{S}$ such that,
-$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}.$
+$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}.$
+Now let’s choose $(p^{0},\lambda^{0})\in[S]^{(a)-(e)}$ such that,
+$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}.$
 Then,
-$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}$
-and thus, $(p^{0},\lambda^{0})\in X_{S,f}$ by the definition of $X_{S,f}$. Hence, $(p^{0},\lambda^{0})\in X^{\prime}_{S}\cap X_{S,f}$
+$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|},$
+and thus, $(p^{0},\lambda^{0})\in[S]_{f}$ by the definition of $S]_{f}$. Hence, $(p^{0},\lambda^{0})\in[S]^{(a)-(e)}\cap[S]_{f}$. ∎
 2.2 Construction of a finite set containing an extremal projection cycle-vector
-Above we study some structural properties of optimal projection cycles on certain finite sets.
-Below, we will use this properties to prove that there is a projection cycle on an explicitly given finite set that gives the supremum (2.4).
+Above, we study some structural properties of optimal projection cycles on certain finite sets.
+Below, we will use these properties to prove that there is a projection cycle on an explicitly given finite set that gives the supremum (2.3).
+Let $f:\Omega\rightarrow\mathbb{R}$ be a given continuous function of three variables satisfying conditions (2.1).
 We start by defining the desired set. Let’s name the corners of the unit cube from $T_{1}$ to $T_{8}$ as
@@
 $g_{7}(x,y,z)=f(x,y,z)-f(0,1,z)-f(0,y,1)-f(x,0,1),$
-$g_{8}(x,y,z)=f(x,y,z)-f(0,y,1)-f(x,0,1)-f(1,0,z),$
+$g_{8}(x,y,z)=f(x,y,z)-f(0,y,1)-f(x,0,1)-f(1,0,z)$
 on the unit cube and
@@
 $h_{5}(0,y,z)=f(0,y,z)-f(1,0,z)-f(1,y,0),$
-$h_{6}(1,y,z)=f(1,y,z)-f(0,1,z)-f(0,y,1).$
+$h_{6}(1,y,z)=f(1,y,z)-f(0,1,z)-f(0,y,1)$
 on the faces of the unit cube lying on the planes $z=0,z=1,y=0,y=1,x=0$ and $x=1$ correspondingly.
-Let $M_{i},1\leq i\leq 8$ be a maximum point of the function $g_{i}$ on the domain $\Omega=[0,1]^{3}$. And $F_{i},1\leq i\leq 6$ be a maximum point of the function $h_{i}$ on the face it’s defined.
+Let $M_{i},1\leq i\leq 8$ be a maximum point of the function $g_{i}$ on the domain $\Omega=[0,1]^{3}$. And $F_{i},1\leq i\leq 6$ be a maximum point of the function $h_{i}$ on the face it is defined.
 Let
-$U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$
+$U=U(f)=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\},$
 and let $U_{x},U_{y},U_{z}$ be the set of projections of elements of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively.
-We will see that, the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ will not change the correctness of the following theorem.
+We will see that the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ will not change the correctness of the following theorem.
 Remark 2.1.
-[…]
+Some of the points $T_{i},\ 1\leq i\leq 8$, $M_{j},\ 1\leq j\leq 8$, $F_{k},\ 1\leq k\leq 6$ may coincide. In this case, the set $U$ will narrow.
 Theorem 2.2.
 Let $f\in C(\Omega)$ be a function satisfying the inequalities (2.1) and $S=S_{x}\times S_{y}\times S_{z}\subset\Omega$ be any arbitrary finite set with $U\subset S$, then we have:
-$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$ (2.12)
+$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.11)
 Proof.
-Let, $X^{\prime\prime}_{S}$ be the set of projection cycle-vectors $(p,\lambda)\in X_{S}$ that satisfies the propositions $(b),(c),(d)$ of the Theorem 2.2.
-It is clear that, $X^{\prime}_{S}\subset X^{\prime\prime}_{S}$ and hence from the Lemma 2.5 we have,
-$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
+Let, $S]^{(a)-(d)}$ be the set of projection cycle-vectors $(p,\lambda)\in[S]$ that satisfies propositions $(a),(b),(c),(d)$ of the Theorem 2.2.
+It is clear that $S]^{(a)-(e)}\subset[S]^{(a)-(d)}$, and hence from Lemma 2.5 we have
+$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(d)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
 Thus, it is enough to prove that
-$\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
-Assume the contrary.
-Since, $U\subset S$ and $S=S_{x}\times S_{y}\times S_{z}$ we have $U_{x}\subset S_{x}$, $U_{y}\subset S_{y}$ and $U_{z}\subset S_{z}$ which imply $U_{x}\times U_{y}\times U_{z}\subset S$ and therefore we can write the inequality
-$\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
-Hence, there exist a projection cycle $(p^{0},\lambda^{0})\in X^{\prime\prime}_{S}$ such that
-$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.13)
-Clearly, $(p^{0},\lambda^{0})\not\in X_{U_{x}\times U_{y}\times U_{z}}$.
-If all points in $p^{0}$ with positive weight belong to $U$, then the set of the projections of them into the edges of the unit cube (which are all the possible locations of points of $p^{0}$ with negative weights) would belong to the set $U_{x}\times U_{y}\times U_{z}$ and hence we would have $p^{0}\subset U_{x}\times U_{y}\times U_{z}$.
-Thus, there exist a point $x\in p^{0}$ with positive weight($\lambda^{0}(x)>0$) such that, $x\not\in U$.
-Moreover, let $(p^{0},\lambda^{0})$ be chosen to minimise the number of positive weighted points not belonging to the set $U$ among all projection cycle-vectors $(p,\lambda)\in X^{\prime\prime}_{S}$ satisfying (2.13).
-By and Lemma 2.5 and Theorem 2.2 part $(d)$, we have that $x$ cannot be on the edges of the unit cube. Thus, we have the following cases:
+$\max_{(p,\lambda)\in[S]^{(a)-(d)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
+Since $U\subset S$ and $S=S_{x}\times S_{y}\times S_{z}$ we have $U_{x}\subset S_{x}$, $U_{y}\subset S_{y}$ and $U_{z}\subset S_{z}$ which imply $U_{x}\times U_{y}\times U_{z}\subset S$ and therefore we can write the inequality
+$\max_{(p,\lambda)\in[S]^{(a)-(d)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
+Assume the contrary to the theorem, let there exists a projection cycle $(p^{0},\lambda^{0})\in[S]^{(a)-(d)}$ such that
+$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.12)
+Clearly, $(p^{0},\lambda^{0})\not\in[U_{x}\times U_{y}\times U_{z}]$. Moreover we can choose it to be optimal in the set $S$ for the function $f$, that is, $(p^{0},\lambda^{0})\in[S]_{f}$.
+If all points in $p^{0}$ with positive weight belong to $U$, then the set of their projections on the edges of the unit cube (which are all the possible locations of points of $p^{0}$ with negative weights) would belong to the set $U_{x}\times U_{y}\times U_{z}$, and hence we would have $p^{0}\subset U_{x}\times U_{y}\times U_{z}$.
+[…]
+We define our moves based on the location of $x$ in the unit cube. Thus, we have the following cases:
 $x$ is an interior point of one of the faces of the unit cube $\Omega$.
@@
 Consider the planes $x=\xi$ and $y=\gamma$.
-In each of these planes, at least one point of $(p^{0},\lambda^{0})$ with negative weight should exist.
+In each of these planes, there should be at least one point $(p^{0},\lambda^{0})$ with negative weight.
 By Theorem 2.2, part $(c)$, these negative weighted points could only be on the edges, not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$.
-The intersections of the planes $x=\xi$ and $y=\gamma$ and the edges are the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ respectively.
-Since both of $(\xi,1,0)$ and $(1,\gamma,0)$ are well-ordered with $x$ and lying on an interior plane, by the Theorem 2.2, part $(b)$, none of $(\xi,1,0)$ and $(1,\gamma,0)$ can have a negative weight.
+The intersections of the planes $x=\xi$ and $y=\gamma$ and the edges are the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$, respectively.
+Since both $(\xi,1,0)$ and $(1,\gamma,0)$ are well-ordered with $x$ and lie on an interior plane, by Theorem 2.2, part $(b)$, none of $(\xi,1,0)$ and $(1,\gamma,0)$ can have a negative weight.
 Thus, both $(\xi,0,1)$ and $(0,\gamma,1)$ have negative weights.
@@
 $|\lambda^{0}(\xi,0,1)|\geq\lambda^{0}(x)\text{ and }|\lambda^{0}(0,\gamma,1)|\geq\lambda^{0}(x)$
-and thus, $\min(|\lambda^{0}(\xi,0,1)|,|\lambda^{0}(0,\gamma,1)|,|\lambda^{0}(x)|)=\lambda^{0}(x)$.
-Let’s move the weights on the points $x$,$(\xi,0,1)$ and $(0,\gamma,1)$ to the points $F_{1}=(\xi_{1},\gamma_{1},0)$ and its two corresponding projections, namely $(\xi_{1},0,1)$ and $(0,\gamma_{1},1)$.
-Or in other words, we define
+and thus $\min(|\lambda^{0}(\xi,0,1)|,|\lambda^{0}(0,\gamma,1)|,|\lambda^{0}(x)|)=\lambda^{0}(x)$.
+[…]
 $p^{\prime}=(x,(\xi,0,1)\,(0,\gamma,1),(\xi_{1},\gamma_{1},0),(\xi_{1},0,1),(0,\gamma_{1},1))$
 and vector
-$\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$
-where, $\varepsilon=\lambda^{0}(x)$.
+$\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon),$
+where $\varepsilon=\lambda^{0}(x)$.
 Let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon})$.
@@
 $\varepsilon=\min(|\lambda^{0}(\xi,0,1)|,|\lambda^{0}(0,\gamma,1)|,|\lambda^{0}(x))|,$
-by the Lemma 2.2 we have
-$\sum|\lambda^{1}_{i}|\leq\sum|\lambda^{0}_{i}|$
+by Lemma 2.2 we have
+$\sum|\lambda^{1}_{i}|\leq\sum|\lambda^{0}_{i}|.$
 On the other hand,
 $\sum\lambda^{1}_{i}f(x_{i})-\sum\lambda^{0}_{i}f(x_{i})=$
-$=f(\xi_{1},\gamma_{1},0)-f(\xi_{1},1,0)-f(1,\gamma_{1},0)-f(\xi,\gamma,0)+f(\xi,1,0)+f(1,\gamma,0)=h(\xi_{1},\gamma_{1},0)-h(\xi,\gamma,0)\geq 0$
-where the inequality follows from the choice of $F_{1}=(\xi_{1},\gamma_{1},0)$ as maximum of the function $h_{1}$. Thus, we get
-$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
-Clearly, $(p^{1},\lambda^{1})\in X^{\prime\prime}_{S}$ as no negative weighted point became positive weighted and vice versa.
-Moreover, $\lambda^{1}(x)=\lambda^{0}(x)-\lambda_{\varepsilon}(x)=0$ and hence, $(p^{1},\lambda^{1})$ has less number of positive weighted points not belonging $U$ than $(p^{0},\lambda^{0})$ has.
-Thus, we get a contradiction to the choice of $(p_{0},\lambda_{0})$.
+$f(\xi_{1},\gamma_{1},0)-f(\xi_{1},1,0)-f(1,\gamma_{1},0)-f(\xi,\gamma,0)+f(\xi,1,0)+f(1,\gamma,0)=h(\xi_{1},\gamma_{1},0)-h(\xi,\gamma,0)\geq 0,$
+where the inequality follows from the choice of $F_{1}=(\xi_{1},\gamma_{1},0)$ as the maximum of the function $h_{1}$. Thus, we get
+$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
+This proves that, $(p^{1},\lambda^{1})\in[S]_{f}$.
+This makes the weight at point $x$ equal to zero in the new projection cycle.
+A similar movement of the weights can be constructed if $x$ lies on one of the other faces too.
 Case 2: $x$ is an interior point of the unit cube $\Omega$.
@@
 Consider the planes $x=\xi$, $y=\gamma$, $z=\zeta$.
-Their intersections with the edges of the unit cube that are not adjacent to the vertices not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$ will be the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ and $(1,0,\zeta),(0,1,\zeta)$ respectively.
+Their intersections with the edges of the unit cube that are not adjacent to the vertices not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$ will be the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ and $(1,0,\zeta),(0,1,\zeta)$, respectively.
 To make the sum over the corresponding plane zero, at least one (maybe both) of two points in each plane should belong to $(p^{0},\lambda^{0})$ with a negative weight.
@@
 For each $G_{i}$ we proceed by moving the weights at point $x$ and elements of $G_{i}$ to $M_{i}$ and its corresponding projections on the edges.
-For example, if all elements of $G_{1}$ belong to $p^{0}$, then we construct $p^{\prime}=(x,(\xi,0,1),(1,\gamma,0),(0,1,\zeta),(\xi_{1},\gamma_{1},\zeta_{1}),(\xi_{1},0,1),(1,\gamma_{!},0),(0,1,\zeta_{1}))$ and vector $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon,-\varepsilon)$ where $M_{1}=(\xi_{1},\gamma_{1},\zeta_{1})$ is the maximum point of $g_{1}$ defined above and $\varepsilon>0$.
-As in Case 1, let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon}).$ Since, at each of the first four points of $(p^{\prime},\lambda_{\varepsilon})$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs, by the Lemma 2.2 for $\varepsilon=\min(\lambda)$ we have
+For example, if all elements of $G_{1}$ belong to $p^{0}$, then we construct $p^{\prime}=(x,(\xi,0,1),(1,\gamma,0),(0,1,\zeta),(\xi_{1},\gamma_{1},\zeta_{1}),(\xi_{1},0,1),(1,\gamma_{1},0),(0,1,\zeta_{1}))$ and vector $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon,-\varepsilon)$ where $M_{1}=(\xi_{1},\gamma_{1},\zeta_{1})$ is the maximum point of $g_{1}$ defined above and $\varepsilon>0$.
+As in Case 1, let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon}).$ Since, at each of the first four points of $(p^{\prime},\lambda_{\varepsilon})$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs, by Lemma 2.2 for
+$\varepsilon=\min(|\lambda(x)|,|\lambda((\xi,0,1))|,|\lambda((1,\gamma,0))|,|\lambda((0,1,\zeta))|$
+we have
 $\sum|\lambda_{i}^{1}|\leq\sum|\lambda_{j}^{0}|.$
@@
 Thus, we have
-$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}$
-As in the first case the last inequality contradicts our assumption of $(p^{0},\lambda^{0})$. Similarly for all other sets we will get a contradiction.
-Since all cases lead to the contradiction, the theorem is proved.
-The proof implies that, the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ does not affect to the correctness of the theorem.
-Now we are ready to prove the Theorem $2.1$.
-Proof (Theorem 2.1).
-Applying the Theorem 2.3 in the relation (2.5) we get
-$E(f,\Omega)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{(U_{x}\cup A)\times(U_{y}\cup B)\times(U_{z}\cup C)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=$
-$\sup_{A,B,C}\left(\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\right)=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
+$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}.$
+Again, this proves that, $(p^{1},\lambda^{1})\in[S]_{f}$.
+On the other hand, choice of $\varepsilon$ makes one of the weights of the points $x,(\xi,0,1),(1,\gamma,0),(0,1,\zeta)$ to have zero weight in $(p^{1},\lambda^{1})$.
+Hence, after a finite number of copies of the above change for some sets $G_{i}$, the weight in $x$ will be equal to zero in the resulting projection cycle-vector.
+Thus, we can proceed a finite number of described moves and obtain a projection cycle-vector from $U_{x}\times U_{y}\times U_{z}]$, which is optimal in $S$. ∎
+The proof implies that the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ does not affect the correctness of the theorem. Now, we are ready to state our main result.
+Theorem 2.3.
+Let $f\in C(\Omega)$ be a function satisfying the relations (2.1). Then
+$E(f,\Omega)=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|},$ (2.13)
+where $U_{x},U_{y},U_{z}$ are the projections of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively.
+Proof.
+Applying Theorem 2.2 to the relation (2.4) we get
+$\begin{split}E(f,\Omega)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=\\ =\sup_{A,B,C}\left({\max_{(p,\lambda)\in[(U_{x}\cup A)\times(U_{y}\cup B)\times(U_{z}\cup C)]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=\\ =\sup_{A,B,C}\left(\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\right)=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.\end{split}$ (2.14)
+Hence, we are finished by the Lemma 2.4 and Lemma 2.5. ∎
 Remark 2.2.
-The proofs of the Theorem 2.3 and the Theorem 2.1 shows that, there exists a projection cycle-vector $(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}$ such that $(p,\lambda)$ satisfies all the propositions of the Theorem 2.2 and
+The above-mentioned papers [11],[12], [20] give similar theorems.
+Their set corresponding to $U$ in our case, was the corners of the unit square in papers [11] and [12], and points from a predefined grid in [20].
+While all these are independent of $f$, $U$ depends on the function $f$ in the above theorem.
+Remark 2.3.
+The Theorem 2.3 shows that there exists a minimal projection cycle-vector $(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]$ such that $(p,\lambda)$ satisfies all the propositions of Theorem 2.2 and
 $E(f,\Omega)=\frac{\sum_{x_{i}\in p}\lambda(x_{i})f(x_{i})}{\sum_{x_{i}\in p}|\lambda(x_{i})|}.$
-Moreover, it can be reconstructed by only knowing the weights in the vertices and all positive weighted points.
+In the Appendix, we have shown that there are at most 123 such projection cycle-vectors in the set $U$.
+Moreover, any of them can be reconstructed only by knowing the weights in the vertices and all the positive weighted points(if two of those points coincide, then thinking them as separate and adding their weights after the weight calculations are done) by using the definitions of $M_{i}$ and $F_{i}$.
+For example, if $M_{i}$ has weight $\lambda$, we will add $\lambda$ to all of the points on $G_{i}$.
 3 Minimal Projection Cycles on a Finite Set
-The relation (2.2) implies that, the error can be found by calculating a maximum over a finite set of the form $V=U_{x}\times U_{y}\times U_{z}$.
-In this section we will search for $X^{\prime}_{V,min}$.
 Let $p=(x_{1},x_{2},...,x_{k})$ be an ordered finite subset of $\Omega$ with $k$ points.
@@
 $A\lambda=0$ (3.1)
-for a matrix $A$. Below, we will prove some necessary conditions for an ordered subset $p^{\prime}$ of $p$ to be an optimal projection cycle.
-We start with a well-known algebraic definition and a lemma (See: Definition 1 and Proposition 7 in [22]).
+for a matrix $A$. Below, we will prove some necessary conditions for an ordered subset $p^{\prime}$ of $p$ to be a minimal projection cycle.
+We start with a well-known algebraic definition and a lemma (see: [37, Definition 1 and Proposition 7]).
 Definition 3.1.
@@
 and call it the support of $v$.
-A solution $v_{0}$ of the homogeneous linear equation $Mv=0$ is called a minimal solution if the set $\operatorname{supp}(v_{0})$ is minimal, that is, there exist no other solution $v$ of the equation satisfying
-$\operatorname{supp}(v)\varsubsetneqq\operatorname{supp}(v_{0})$
+A solution $v_{0}$ of the homogeneous linear equation $Mv=0$ is called a minimal solution if the set $\operatorname{supp}(v_{0})$ is minimal, that is, there exists no other solution $v$ of the equation satisfying
+$\operatorname{supp}(v)\varsubsetneqq\operatorname{supp}(v_{0}).$
 Lemma 3.1.
-Let $M$ be a matrix then a solution $v_{0}$ of the equation $Mv=0$ which $i_{1},i_{2},...,i_{t}$-th components are nonzero and all other components are zero, is minimal, then the column system consisting $i_{1},i_{2},...,i_{k}$-th columns of the matrix $M$ is a minimal linearly dependent column system.
+Let $M$ be a matrix then a solution $v_{0}$ of the equation $Mv=0,$ which $i_{1},i_{2},...,i_{t}$ -th components are nonzero and all other components are zero, is minimal, then the column system consisting of $i_{1},i_{2},...,i_{k}$ -th columns of the matrix $M$ is a minimal linearly dependent column system.
+The next lemma connects minimal solutions and minimal projection cycles.
 Lemma 3.2.
-Let $p^{\prime}=(x_{i_{1}},x_{i_{2}},\dots x_{i_{t}})\subset p$ be a minimal projection cycle and $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$.
-Then the vector, $\lambda=(\lambda_{1},...,\lambda_{k})$ is a minimal solution of the equation (3.1), where,
-$\lambda_{i}=\begin{cases}\lambda^{\prime}_{i},\hskip 13.0pt\text{ if }i\in\{i_{1},i_{2},\dots,i_{t}\}\\ 0,\hskip 17.0pt\text{ otherwise }\end{cases}$ (3.2)
+Let $p^{\prime}=(x_{i_{1}},x_{i_{2}},\dots x_{i_{t}})\subset p$ be a minimal projection cycle, and $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$.
+Then the vector, $\lambda=(\lambda_{1},...,\lambda_{k})$ is a minimal solution of the equation (3.1), where
+$\lambda_{i}=\begin{cases}\lambda^{\prime}_{i},\hskip 13.0pt\text{ if }i\in\{i_{1},i_{2},\dots,i_{t}\}\\ 0,\hskip 17.0pt\text{ otherwise.}\end{cases}$ (3.2)
 Proof.
-Assume the contrary, suppose $\lambda$ is not a minimal solution of (3.1). Then, there exist a solution $\lambda^{0}$ of (3.1) such that,
+Assume the contrary; suppose that $\lambda$ is not a minimal solution of (3.1). Then, there exists a solution $\lambda^{0}$ of (3.1) such that
 $\operatorname{supp}(\lambda^{0})\varsubsetneqq\operatorname{supp}(\lambda).$
-Thus, there exist a nonempty set $\{j_{1},j_{2},...,j_{s}\}\varsubsetneqq\{1,2,...,t\}$ such that the $i_{j_{1}},i_{j_{2}},...,i_{j_{s}}$-th components of $\lambda^{0}$ are nonzero and all other components are zero.
+Thus, there exists a nonempty set $\{j_{1},j_{2},...,j_{s}\}\varsubsetneqq\{1,2,...,t\}$ such that the $i_{j_{1}},i_{j_{2}},...,i_{j_{s}}$-th components of $\lambda^{0}$ are nonzero and all other components are zero.
 And, this implies that
@@
 is a projection cycle with vector $\lambda^{\prime\prime}=(\lambda^{0}_{i_{j_{1}}},\lambda^{0}_{i_{j_{2}}},...,\lambda^{0}_{i_{j_{s}}})$, which contradicts the minimality of $p^{\prime}$. ∎
-The next theorem gives a necessary condition on minimality of a projection cycle.
+The next theorem gives a necessary condition on the minimality of a projection cycle.
 Theorem 3.1.
@@
 Proof.
-Let $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$ and the vector $\lambda$ be defined as in (3.2).
-Then, by the Lemma 3.2, $\lambda$ is a minimal solution of (3.1).
-Thus, by the Lemma 3.1 the proof follows.
+Let $\lambda^{\prime}=(\lambda^{\prime}_{i_{1}},\lambda^{\prime}_{i_{2}},\dots\lambda^{\prime}_{i_{t}})$ be a vector assigned to $p^{\prime}$ and let the vector $\lambda$ be defined as in (3.2).
+Then, according to Lemma 3.2, $\lambda$ is a minimal solution of (3.1).
+Thus, by Lemma 3.1 the proof follows.
 ∎
-We can construct a system of equation with variables being the weights at points of $U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$ and equations being equations on the faces $x=0$, $x=1$, $y=0$, $y=1$, $z=0$ and $z=1$ correspondingly, as all other variables in (3.1) can be defined uniquely by the Remark 2.2.
-It is easy to verify that, this new system will be
+Appendix: Finding all minimal projection cycles
+The relation (2.14) implies that the error can be found by calculating a maximum over a finite set of the form $V=U_{x}\times U_{y}\times U_{z}$.
+Below, we will search for $V]^{(a)-(e)}_{min}-$the set of all minimal projection cycle-vectors on set $V$ satisfying all the propositions from the Theorem 2.1.
+We can construct a system of equations with variables being the weights at the points of $U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$ and equations being equations on the faces $x=0$, $x=1$, $y=0$, $y=1$, $z=0$ and $z=1$ correspondingly, since the weights at all other points in $V$ can be defined uniquely as described in Remark 2.3.
+It is easy to verify that this system will be
 $Mv=0,$
@@
 and $v=(\lambda(T_{1}),\lambda(T_{8}),\lambda(M_{1}),\lambda(M_{2}),\dots,\lambda(M_{8}),\lambda(T_{2}),\lambda(T_{3}),\dots,\lambda(T_{7}),\lambda(F_{1}),\lambda(F_{2}),\dots,\lambda(F_{6}))^{\mathrm{T}}.$
-Reducing the matrix $M$ to row echelon form by using the fact that the rows $r_{1},\dots,r_{6}$ of the matrix satisfies the identity $r_{1}+r_{2}=r_{3}+r_{4}=r_{5}+r_{6},$ gives $M\sim M^{\prime}$, where $M^{\prime}$ is the matrix given below
-$\begin{pmatrix}1&0&-1&-1&-1&-2&-2&-1&0&0&1&1&1&0&0&0&1&0&-1&-2&-1&0\\ 0&1&-1&-1&0&-1&-2&-2&-1&0&0&1&1&1&0&0&0&1&0&-1&-2&-1\\ 0&0&0&0&1&2&1&-1&-2&-1&-1&-1&0&1&1&0&-2&0&2&2&0&-2\\ 0&0&0&0&-1&1&2&1&-1&-2&0&-1&-1&0&1&1&-2&-2&0&2&2&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\end{pmatrix}$
-By the Lemma 2.5 we know that there exist an optimal projection cycle on $U_{x}\times U_{y}\times U_{z}$ that has $\lambda(T_{1})>0$, $\lambda(T_{8})>0$, $\lambda(M_{i})\geq 0$, $1\leq i\leq 8$, $\lambda(F_{i})\geq 0$, $1\leq i\leq 6$, $\lambda(T_{i})\leq 0$, $2\leq i\leq 7$ and $\lambda(F_{i})\lambda(F_{j})=0$ if $i\neq j$ and $j-i$ is even.
-Thus, its enough to find all minimal solutions of $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0.$
-Assume $v$ is a minimal solution of the equation $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0$.
-The Theorem 3.1 implies that the system of columns of $M^{\prime}$ corresponding to the nonzero components of $v$ is a minimal linear dependent system.
-It is not hard to verify that there are $123$ minimal column systems containing the first two columns and satisfying the above conditions.
-And these are the ones given below.
+Reducing the matrix $M$ to the row echelon form by using the fact that the rows $r_{1},\dots,r_{6}$ of the matrix satisfy the identity $r_{1}+r_{2}=r_{3}+r_{4}=r_{5}+r_{6},$ gives $M\sim M^{\prime}$, where $M^{\prime}$ is the matrix given below
+[…]
+By the Lemma 2.5 we know that there exists an optimal projection cycle on $U_{x}\times U_{y}\times U_{z}$ that has $\lambda(T_{1})>0$, $\lambda(T_{8})>0$, $\lambda(M_{i})\geq 0$, $1\leq i\leq 8$, $\lambda(F_{i})\geq 0$, $1\leq i\leq 6$, $\lambda(T_{i})\leq 0$, $2\leq i\leq 7$ and $\lambda(F_{i})\lambda(F_{j})=0$ if $i\neq j$ and jij-i is even.
+Thus, it is sufficient to find the maximum over all minimal projection cycles on $V$ satisfying $\lambda(T_{1}),\lambda(T_{8})>0.$ Thus, by the Theorem 3.1, we can find all such projection cycles by finding all minimal solutions of the system $M^{\prime}v=0$ with $\lambda(T_{1}),\lambda(T_{8})>0$ or in other words, finding all minimal linearly dependent column systems containing the first two columns.
+It is not hard to verify with the help of a computer that there are $123$ minimal column systems containing the first two columns and corresponding to a minimal solution satisfying the above conditions.
+And corresponding projection cycle-vectors are the ones given below.
 Note that we give only the weights on the corner points and the points with positive weights.
@@
 for the minimal projection cycle
-1.1. $T_{1},T_{8},T_{2},T_{5}:\ \lambda(T_{1})=1,\ \lambda(T_{8})=1,\ \lambda(T_{2})=-1,\ \lambda(T_{5})=-1$
+$1.$ $T_{1},T_{8},T_{2},T_{5}:\ \lambda(T_{1})=1,\ \lambda(T_{8})=1,\ \lambda(T_{2})=-1,\ \lambda(T_{5})=-1$
 this is the function $f(x,y,z)=xz$:
@@
 2.11. $T_{1},T_{8},M_{5},T_{3},T_{7}:\quad\lambda(T_{1})=2,\quad\lambda(T_{8})=3,\quad\lambda(M_{5})=1,\quad\lambda(T_{3})=-2,\quad\lambda(T_{7})=-1$
-[…]
+this is the function $f(x,y,z)=\begin{cases}xz,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ for\ \ 0\leq x\leq\frac{1}{2},\ 0\leq z\leq\frac{1}{2},\\ \frac{z}{2}+\frac{(2x-1)y}{4},\ \ \ \ \ \ \ \ \ \ \ \ \ for\ \ \frac{1}{2}\leq x\leq 1,\ 0\leq z\leq\frac{1}{2},\\ \frac{x}{2}+\frac{(2z-1)y}{4},\ \ \ \ \ \ \ \ \ \ \ \ \ for\ \ 0\leq x\leq\frac{1}{2},\ \frac{1}{2}\leq z\leq 1,\\ \frac{1}{4}+\frac{(2x-1)y}{4}+\frac{(2z-1)y}{4},\ for\ \ \frac{1}{2}\leq x\leq 1,\ \frac{1}{2}\leq z\leq 1\end{cases}$:
 $E(f,\Omega)=\frac{7}{48}=\frac{\sum_{i=1,8,5,7}\lambda(T_{i})f(T_{i})+\lambda(M_{5})f(M_{5})+\sum_{i=3,4,5}\lambda(M_{5,l_{i}})f(M_{5,l_{i}})}{\sum_{i=1,8,5,7}|\lambda(T_{i})|+|\lambda(M_{5})|+\sum_{i=3,4,5}|\lambda(M_{5,l_{i}})|},$
-where $M_{5}=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, $M_{5,l_{i}}=\Pr_{l_{i}}M_{5}$, $i=3,4,5$, etc.
+where $M_{5}=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, $M_{5,l_{i}}=\Pr_{l_{i}}M_{5}$, $i=3,4,5$.